### v1

# Can LLMs handle WebShell detection? Overcoming Detection Challenges with Behavioral Function-Aware Framework

Abstract
WebShell attacks, in which malicious scripts are injected into web servers, are a major cybersecurity threat. Traditional machine learning and deep learning methods are hampered by issues such as the need for extensive training data, catastrophic forgetting, and poor generalization. Recently, Large Language Models (LLMs) have gained attention for code-related tasks, but their potential in WebShell detection remains underexplored. In this paper, we make two major contributions: (1) a comprehensive evaluation of seven LLMs, including GPT-4, LLaMA 3.1 70B, and Qwen 2.5 variants, benchmarked against traditional sequence- and graph-based methods using a dataset of 26.59K PHP scripts, and (2) the Behavioral Function-Aware Detection (BFAD) framework, designed to address the specific challenges of applying LLMs to this domain. Our framework integrates three components: a Critical Function Filter that isolates malicious PHP function calls, a Context-Aware Code Extraction strategy that captures the most behaviorally indicative code segments, and Weighted Behavioral Function Profiling (WBFP) that enhances in-context learning by prioritizing the most relevant demonstrations based on discriminative function-level profiles. Our results show that larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit the opposite trade-off. However, all models lag behind previous State-Of-The-Art (SOTA) methods. With BFAD, the performance of all LLMs improved, with an average F1 score increase of 13.82%. Larger models such as GPT-4, LLaMA 3.1 70B, and Qwen 2.5 14B outperform SOTA methods, while smaller models such as Qwen 2.5 3B achieve performance competitive with traditional methods. This work is the first to explore the feasibility and limitations of LLMs for WebShell detection, and provides solutions to address the challenges in this task.
1 Introduction
The rapid growth of web applications and the expansion of cloud-based services have escalated the cybersecurity threat landscape, with WebShells emerging as a significant concern. WebShells are malicious scripts injected into web servers that allow attackers to remotely execute arbitrary commands, steal sensitive data, and compromise system integrity (Starov et al., 2016; Cui et al., 2018). According to a Cisco Talos report (Talos, 2024), threat groups used a variety of web shells against vulnerable or unpatched web applications in 35% of incidents in Q4 2024, a sharp increase from the previous quarter, when such activity was observed in only 10% of incidents. These attacks are particularly dangerous due to the stealthy nature of WebShells, which are rapidly evolving to evade traditional detection methods (Hannousse & Yahiouche, 2021).
In response to this challenge, the community has proposed several approaches. Rule-based methods that rely on predefined signatures or heuristics are increasingly ineffective against the complexity and diversity of modern WebShells (Le et al., 2021; Jinping et al., 2020). Machine learning models, especially deep learning techniques (Pu et al., 2022), have shown promise in addressing these threats. However, they require extensive training on large datasets - resources that are often difficult to obtain and sensitive in nature  (Shang et al., 2024). In addition, these models face challenges such as catastrophic forgetting and poor generalization, especially when dealing with obfuscated or encrypted attacks  (Jinping et al., 2020; Zhang et al., 2025).
Recently, Large Language Models (LLMs) have gained attention for a variety of code-related tasks, including code generation  (Ma et al., 2024) and vulnerability detection  (Liu & He, 2023; Wang et al., 2025). Studies have shown that with timely engineering, LLMs can perform remarkably well without additional training  (Nong et al., 2024; Trad & Chehab, 2025). They also provide interpretable explanations for their decisions, which is critical in cybersecurity contexts (Ma et al., 2024). Despite these advantages, there has been limited exploration of LLMs for WebShell detection.
Detecting WebShells with LLMs presents unique challenges that differ from other code analysis tasks. WebShells often employ obfuscation and encryption techniques, and are embedded in large codebases dominated by benign content (Liu & He, 2023). For example, the largest WebShell in our dataset spans 1,386,438 tokens, far exceeding the context window of most LLMs, which risks truncating critical malicious segments when processing entire source files (Wang et al., 2025; Ceka et al., 2024). In addition, in-context learning (ICL) struggles in this domain: the variability and obfuscation of WebShells complicates the selection of effective demonstrations, and these examples further occupy significant context space, reducing capacity for the target code (Yuan et al., 2024). While recent research has focused on increasing the context length of LLMs (Chen et al., 2023), studies suggest that the performance of an LLM tends to degrade with longer inputs, and the low processing speed may also become unacceptable for practical use (Ma et al., 2024; Fang et al., 2024).
In this paper, we present two key contributions to advance the application of LLMs in WebShell detection.
First, we systematically evaluate LLMs in the context of WebShell detection, comparing their performance with traditional state-of-the-art machine learning (SOTA) methods. Specifically, we evaluate seven closed-source and open-source LLMs of different sizes, including GPT-4 (Achiam et al., 2023), LLaMA 3.1 70B (Grattafiori et al., 2024), Qwen 2.5 Coder (14B/3B) (Yang et al., 2024), and Qwen 2.5 (3B/1.5B/0.5B) (Yang et al., 2024) on a dataset containing 26.59K PHP scripts (4.93K WebShells and 21.66K benign samples). Our analysis reveals several key findings:
  • Larger LLMs, such as GPT-4 and Qwen 2.5 Coder 14B, achieve near perfect precision (close to 100%), but struggle with recall (e.g., GPT-4’s recall is 85.98%), indicating difficulty in detecting all malicious instances.
  • Smaller LLMs, such as Qwen 2.5 Coder 3B and Qwen 2.5 0.5B, have high recall (close to 100%) but suffer from low precision (e.g. Qwen 2.5 Coder 3B’s precision is 38.93%), indicating a tendency to misclassify benign files as malicious.
  • Randomly selected ICL demonstrations degrade LLM detection performance; examples selected based on semantic similarity to the source code do not yield significant improvements.
We compare the performance of LLMs with traditional methods, including Glove+SVM (Petridis, 2024; Rigutini et al., 2024), CodeBERT+Random Forest (Alghamdi et al., 2022), and graph-based approaches such as GAT (Kang et al., 2023). The best performing LLM, Qwen 2.5 Coder 14B, achieves an F1 score of 96.39%, although it still lags behind GAT-based methods, which achieve an F1 score of 98.87%. [^1] While traditional models such as GAT have demonstrated superior performance, they require extensive training and significant data collection efforts. In contrast, LLMs enable direct detection through prompts, leveraging their pre-trained knowledge with minimal task-specific resources to deliver competitive results..
Second, we present the Behavioral Function-Aware Detection (BFAD) framework to address the identified challenges and to improve the performance of LLMs to meet the requirements of downstream applications by achieving a more balanced trade-off between precision and recall. Our framework combines risk-filtering techniques with an enhanced ICL strategy that uses weighted demonstration selection to prioritize examples most closely related to key malicious behaviors. Experimental results show that our approach improves the average F1 score across all LLMs by 13.82%, with GPT-4 and Qwen 2.5 0.5B improving by 6.89% and 51.32%, respectively. In several cases, such as GPT-4, LLaMA 3.1 70B, Qwen 2.5 Coder 14B, and Qwen 2.5 Coder 3B, our approach enables performance that is competitive or even superior to traditional methods.
To the best of our knowledge, this is the first work to systematically analyze the feasibility and limitations of applying LLMs to WebShell detection.
Figure 1: Overview of the Behavioral Function-Aware Detection framework for WebShell detection. It consists of three components: (a) Critical Function Filter, which identifies PHP functions associated with malicious behavior; (b) Context-Aware Code Extraction, which isolates critical code regions to overcome LLM context limitations; and (c) Weighted Behavioral Function Profiling, which selects ICL demonstrations using a behavior-weighted similarity score.
2 Behavioral Function-Aware Detection Framework
We present the Behavioral Function-Aware Detection (BFAD) framework, a comprehensive solution designed to improve WebShell detection by identifying critical code segments and improving the quality of ICL demonstration selection. The architecture of BFAD, as shown in Figure 1, consists of three primary components: (a) Critical Function Filter, which is based on malicious behavior patterns and filters critical PHP function calls to identify key functions; (b) Context-Aware Code Extraction, which addresses the limitations of LLMs in handling long input sequences by selectively extracting critical code regions; and (c) Weighted Behavioral Function Profiling, which improves demonstration selection by calculating a weighted similarity score based on function-level profiling.
2.1 Critical Function Filter
WebShells typically rely on specific PHP functions that facilitate malicious actions such as code execution, data exfiltration, or obfuscation. However, these functions are often embedded in complex, obfuscated code, making detection difficult. To address this, we are developing a Critical Function Filter that classifies PHP functions into six different behavioral categories: Program Execution, Code Execution, Callback Functions, Network Communication, Information Gathering, and Obfuscation and Encryption. This taxonomy reflects the different roles these functions play in WebShells.
Specifically, the Program Execution category includes functions such as exec and system, which execute system-level commands and are often exploited in WebShells to execute arbitrary payloads. Similarly, Code Execution functions such as eval and preg_replace interpret input as executable code, allowing attackers to inject arbitrary scripts. The Callback Functions category includes functions such as array_map and register_shutdown_function, which allow dynamic invocation of functions often used to obfuscate malicious code.
In addition, Network Communication functions such as fsockopen and curl_init allow remote communication for data exfiltration and command-and-control operations. Information Gathering functions, such as phpinfo and getenv, are used by attackers to gather system details. Finally, Obfuscation and Encryption functions, such as base64_encode and openssl_encrypt, help disguise or encrypt malicious payloads to avoid detection.
Our statistical analysis (detailed in Appendix B) shows that WebShell files use critical functions far more often than benign files. On average, WebShells contain 22.76 calls to critical functions, compared to only 0.74 in benign files, underscoring their behavioral complexity and distinctiveness. [^2] While this disparity in critical function usage is significant, it alone does not reliably distinguish WebShells from benign files, since legitimate scripts may also call these functions. Therefore, we use LLMs for deeper contextual analysis to improve detection accuracy.
2.2 Context-Aware Code Extraction
Building on the Critical Function Filter, we introduce a Context-Aware Code Extraction strategy that identifies and extracts the critical code regions that indicate malicious behavior. These regions focus on the identified critical functions and their surrounding contexts, ensuring that the LLM focuses on the most relevant parts of the code.
The complete extraction procedure is formalized in Algorithm 1, which takes as input the source code $\mathcal{C}$, the list of critical functions $\mathcal{F}$, and the context window size $\tau$, and produces a set of extracted critical code regions $\mathcal{R}$.
We reduce the input size by selectively extracting critical regions of code and merging overlapping segments while preserving behavioral specificity. However, this approach may inadvertently exclude certain global contextual information and increase the emphasis on critical functions, potentially leading to false positives when analyzing benign files that legitimately use such functions. To mitigate this, we append truncated, non-overlapping code segments when context length allows, ensuring that the model receives a balanced representation of local and global code context.
Algorithm 1 Context-Aware Code Extraction
1:  Input: Source code $\mathcal{C}$, list of critical functions $\mathcal{F}$, context window size $\tau$
2:  Output: Extracted critical code regions $\mathcal{R}$
3:  Initialize empty set of regions: $\mathcal{R}\leftarrow\emptyset$
4:  for each function $f\in\mathcal{F}$ do
5:     Locate all occurrences of $f$ in $\mathcal{C}$
6:     for each occurrence of $f$ at position $p$ do
7:        Extract context window $[p-\tau,p+\tau]$ from $\mathcal{C}$
8:        Add the extracted region to $\mathcal{R}$
9:     end for
10:  end for
11:  Merge overlapping regions in $\mathcal{R}$
12:  Compute remaining context budget $B$
13:  if $B>0$ then
14:     Select additional non-overlapping code segments from $\mathcal{C}$
15:     Add selected segments to $\mathcal{R}$
16:  end if
17:  return $\mathcal{R}$
2.3 Weighted Behavioral Function Profiling
Based on the extracted critical code regions, we propose Weighted Behavioral Function Profiling (WBFP), a method that computes a weighted similarity score to identify behaviorally similar examples for ICL effectively. WBFP assigns weights to each function type based on its prevalence and usage in WebShell versus benign files, quantified by three metrics: coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$). The coverage difference measures the proportion of files containing a specific function across the two datasets. The frequency ratio is the ratio of the average number of occurrences of the function per file in WebShell files to the average number per file in benign files. The usage ratio reflects the total number of function occurrences in WebShell files compared to those in benign files. These metrics are combined to calculate a discrimination score for each function type
$\text{Score}_{f}=(r_{c}\cdot\alpha)+(r_{f}\cdot\beta)+(r_{u}\cdot\gamma),$
where $\alpha$, $\beta$, and $\gamma$ are empirically determined weights, set to 1 for balanced contribution in our experiments (see Section 3.2). We normalize the discrimination scores to weights:
$w_{f}=\frac{\text{Score}_{f}}{\sum_{f^{\prime}\in\mathcal{F}}\text{Score}_{f^{\prime}}}.$
WBFP then uses the embeddings $E(\cdot)$ generated by st-codesearch-distilroberta-base (Abi Akl, 2023; Al-Kaswan et al., 2023) to compute the similarity between the files $x$ and $y$. Let $\mathcal{F}$ denote all critical function types. For each function type $f\in\mathcal{F}$, we concatenate critical regions $R_{f}(x)$ and compute their embeddings:
$\mathbf{e}_{f}(x)=E(\text{concat}_{f}(x)),$
The semantic similarity between the files $x$ and $y$ for the function type $f$ is given by
$s_{f}(x,y)=\frac{\mathbf{e}_{f}(x)\cdot\mathbf{e}_{f}(y)}{\|\mathbf{e}_{f}(x)\|\|\mathbf{e}_{f}(y)\|}.$
The final similarity between the files is the weighted sum of the similarities:
$\text{Sim}(x,y)=\sum_{f\in\mathcal{F}}w_{f}\cdot s_{f}(x,y).$
This weighted similarity prioritizes function types critical to WebShell behavior, reducing the impact of irrelevant semantic features. Using this score, WBFP ensures that ICL demonstrations accurately capture malicious patterns, improving detection performance.
2.4 LLM-Based Detection Framework
We integrate the BFAD framework into the LLM-based detection system, which combines the context-aware code extraction strategy and WBFP to optimize the use of LLMs for WebShell detection. By incorporating both critical code regions and global context, along with behaviorally relevant demonstrations, our framework enhances the ability of the LLM to accurately identify malicious code patterns.
The input to the LLM consists of two main components: (a) a system directive that defines the model’s role as a cybersecurity expert, and (b) a user query that contains the extracted critical code segments and a selected ICL demonstration. To balance efficiency and performance, we limit the user query to one ICL demonstration, which reduces computational overhead while preserving sufficient context for reliable detection. Our prompt is described in detail in Appendix A.
3 Experiment
3.1 Dataset Overview
We constructed a comprehensive dataset consisting of 21,665 benign PHP programs and 4,929 webshells. The benign programs were obtained from established open-source PHP projects to ensure applicability to real-world scenarios. The webshells were collected from public security repositories and augmented with synthetic obfuscation techniques to increase diversity.  [^3] We acknowledge the potential for data leakage due to the pre-training of LLMs. However, the autoregressive nature of this process does not explicitly capture WebShell classification, although it may affect other code generation tasks. To further minimize any risk of leakage, we restricted our selection of PHP programs to projects updated between October 2024 and 2025, thus reducing overlap with LLM training corpora. Using GPT-4’s tokenizer, we analyzed the token lengths of both sample types. The webshell samples had a maximum token length of 1,386,438 and an average of 30,856.60 tokens, compared to a maximum of 305,670 tokens and an average of 2,242.89 tokens for benign programs. These results indicate that webshells are typically significantly longer than benign samples. A detailed summary of the dataset composition can be found in the Table 2 in the Appendix C.
3.2 Experiment Setup
ICL Settings.
We randomly selected 60% of the dataset to create a fixed known demonstration library for ICL. Using this subset, we computed normalized scores for different function categories based on the WBFP method, giving equal weight to coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$) to profile functions according to their behavioral importance in distinguishing webshells from benign programs. These scores, detailed in the Table 3 in the Appendix C, guided the selection of ICL demonstrations from the known sample library.
Baseline Models, Hyperparameter Settings, and Evaluation Metrics.
We compared our approach to several baselines: GloVe + SVM, CodeBERT + Random Forest, GCN, and GAT. For GloVe + SVM, we used pre-trained GloVe embeddings with a dimensionality of 300 and an SVM classifier with default parameters (Qi et al., 2018; ZENG et al., 2025). For CodeBERT + Random Forest, we used CodeBERT embeddings with a hidden dimension of 768 and a Random Forest classifier with default settings (Wang et al., 2024a). The GCN was trained with a learning rate of 0.001 over 120 epochs, with 3 hidden layers and a hidden dimension of 32 (Feng et al., 2024). The GAT was trained with a learning rate of 0.001 over 120 epochs, with 3 hidden layers, a hidden dimension of 8, and 8 attention heads (Feng et al., 2024). We evaluated the models using standard classification metrics: accuracy, precision, recall, and F1 score.
4 Results and Analysis
In this section, we systematically evaluate the performance of LLMs of different scales for WebShell detection, and assess the improvements provided by our proposed BFAD framework. Our analysis addresses three research questions (RQs) to explore both the baseline LLM capabilities and the effectiveness of the BFAD components:
  • RQ1: How do large and small scale LLMs perform in WebShell detection compared to traditional ML and DL methods, and how does BFAD improve their effectiveness?
  • RQ2: How effective is context-aware code extraction at balancing global context and local behavioral focus under LLM context length constraints?
  • RQ3: How does WBFP improve demo selection for ICL?
4.1 Performance Evaluation of LLMs and BFAD Enhancements (RQ1)
We evaluated seven LLMs, including large-scale models such as GPT-4, LLaMA-3.1-70B, and Qwen-2.5-Coder-14B, and small-scale models such as Qwen-2.5-Coder-3B, Qwen-2.5-3B, Qwen-2.5-1.5B, and Qwen-2.5-0.5B. These were compared with traditional ML and DL baselines, including sequence-based methods (Glove+SVM, CodeBERT+RF) and graph-based methods (GCN, GAT). The results are detailed in Table 4 in Appendix D.
Characteristics of LLMs in WebShell Detection
Our evaluation reveals distinct performance characteristics of LLMs in WebShell detection that are influenced by their size. Compared to traditional machine learning and deep learning methods, especially GAT, vanilla LLMs exhibit unbalanced performance - large models lack sufficient recall, while small models lack precision - making them less effective without targeted improvements.
In particular, large LLMs such as GPT-4 and Qwen-2.5-Coder-14B achieve high precision (100% and 99.32%, respectively) but moderate recall (85.98% and 93.63%, respectively), reflecting a conservative classification bias. This is likely due to their extensive training on diverse datasets, which prioritizes accurate identification of benign code over detection of rare or novel WebShells. Code-specific fine-tuning improves the performance of Qwen-2.5-Coder-14B over general-purpose models such as LLaMA-3.1-70B (97.31% precision, 92.36% recall), underscoring the benefits of domain specialization. In contrast, small-scale LLMs such as Qwen-2.5-0.5B and Qwen-2.5-1.5B exhibit nearly perfect recall (100% and 95.77%, respectively) but extremely low precision (18.65% and 34.61%, respectively). This imbalance may be due to their increased sensitivity to prompts, making them more likely to provide answers that match user expectations, a problem exacerbated by their limited ability to model complex code relationships.
Effects of BFAD on LLMs
The BFAD framework significantly improves the performance of LLMs for WebShell detection. For large-scale LLMs, BFAD increases recall while maintaining near-perfect precision; in particular, GPT-4’s recall increases by 12.73% to 98.71%, yielding an F1 score of 99.35%, which exceeds the graph-based GAT baseline (98.87% F1), while LLaMA-3.1-70B achieves an F1 score of 98.40%, closely matching GAT. Conversely, for small LLMs, BFAD significantly increases precision without compromising high recall; Qwen-2.5-0.5B’s precision increases by 52.36% to 71.10%, resulting in an F1 score of 82.67% (a 51.23% improvement), and Qwen-2.5-1.5B achieves an F1 score of 65.33% (up 14.85%). These advances enable BFAD-enhanced LLMs to match or exceed traditional methods without the need for additional fine-tuning or training, driven by a strategic focus on critical code regions and the use of weighted behavioral profiling to select contextually relevant demonstrations. This approach effectively mitigates the inherent context length constraints of LLMs and improves their generalization from limited examples. Notably, small-scale LLMs exhibit more pronounced gains, underscoring the critical role of structured input for models with limited capacity, while large-scale LLMs benefit from improved recall, making them suitable for safety-critical applications. These results highlight the transformative potential of integrating domain-specific strategies with LLMs to address specialized challenges, such as WebShell detection, with exceptional effectiveness.
4.2 The Effectiveness of Context-Aware Code Extraction (RQ2)
We evaluated the effectiveness of Context-Aware Code Extraction using two models: GPT-4, a large-scale model, and Qwen-2.5-3B, a smaller-scale model. Three configurations were compared: (1) predictions based on the full source code, (2) predictions using only extracted critical regions, and (3) a hybrid approach combining critical regions with truncated source code. Results are reported in Tables 5 and 6 in Appendix D.
Impact of Critical Regions
For the smaller model, Qwen-2.5-3B, critical regions significantly improve performance over the full source baseline. At $\tau=100$, the F1 score increases from 84.37% to 90.91% (+6.54%), with precision increasing from 78.03% to 86.71% (+8.68%) and recall increasing from 91.84% to 95.54% (+3.70%). This improvement validates the focus on behaviorally relevant code snippets, which reduces irrelevant context and sharpen the focus of the model. However, as the context length ($\tau$) increases, performance decreases - F1 drops to 88.17% at $\tau=300$, likely due to the model’s limited ability to handle extended context, which introduces noise that degrades accuracy.
For the larger model, GPT-4, critical regions increase recall but slightly decrease precision. At $\tau=300$, recall improves from 85.98% to 96.18% (+10.20%), while precision drops from 100.00% to 98.69% (-1.31%). The F1 score increases from 92.46% to 97.42% (+4.96%), indicating that GPT-4 effectively uses localized behavioral cues to improve recall. However, the reduced global context may introduce small biases, leading to a precision trade-off, although the overall performance remains strong due to the model’s greater capacity.
Balancing Precision and Recall with the Hybrid Strategy
The hybrid strategy improves model performance by effectively balancing precision and recall. For Qwen-2.5-3B, it increases precision over using critical regions alone, from 86.71% to 89.02% (+2.31%) for $\tau=100$ and from 82.32% to 85.55% (+3.23%) for $\tau=300$. Although recall decreases slightly, the F1 score increases to 89.70% at $\tau=300$ (+1.53% from 88.17%), indicating that the hybrid approach reduces noise in longer contexts and supports smaller models by maintaining focus on critical regions while integrating valuable global context. For GPT-4, the strategy improves recall without compromising precision: at $\tau=300$, recall increases from 85.98% (using the source code) to 96.82% (+10.84%), while precision remains steady at 100.00%, yielding an F1 score of 98.38%–a 5.92% gain over the source code and a 0.96% improvement over critical regions alone.
4.3 The Effectiveness of WBFP for In-Context Learning (RQ3)
We evaluated the effectiveness of WBFP for ICL demonstration selection using Qwen-2.5-3B and GPT-4. This evaluation builds on the Context-Aware Code Extraction with the hybrid strategy. Five demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt). Results are detailed in Tables 7 and 8 in Appendix D.
Limitations of Random and Semantic Similarity-Based Selection
Random selection makes ICL performance much worse by adding irrelevant examples, which makes both models much less effective. For Qwen-2.5-3B, the F1 score drops to 60.83% under the hybrid strategy with a context length of $\tau=100$, which is a 23.54% reduction from the no-ICL baseline of 84.37%. This is mainly because there is a large drop in precision from 78.03% to 46.33%, although recall remains strong at 88.53%. Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 22.16% decrease from the No-ICL baseline of 98.38%, driven by a drop in precision from 100.00% to 65.58%. These results indicate that Random Selection does not provide contextually relevant demonstrations, making it ineffective for WebShell detection.
The SC-Sim approach, which relies on semantic similarity calculated over entire source code samples, also doesn’t work well. For Qwen-2.5-3B, SC-Sim achieves an F1 score of 84.36%, which is almost the same as the No-ICL baseline (84.37%), with a precision of 78.57% and a recall of 91.08%. For GPT-4, it achieves an F1 score of 96.32%, with perfect precision (100.00%), but a reduced recall of 92.90% compared to the No-ICL baseline’s 96.82%. This limited performance is likely due to the dominance of behaviorally irrelevant code segments in the similarity calculation, which dilutes the focus on critical behavioral patterns essential for accurate WebShell identification.
Superiority of WBFP in Demonstration Selection
For both models, WBFP-Wt consistently outperforms in accuracy, precision, recall, and F1 score, demonstrating its robustness and adaptability for improving ICL in WebShell detection tasks. Specifically, for Qwen-2.5-3B, WBFP-Wt achieves an F1 score of 93.69%, which is 9.33% higher than SC-Sim and 1.28% higher than WBFP-Eq. This is because WBFP-Wt achieves a precision of 88.64% (compared to 78.57% for SC-Sim and 86.39% for WBFP-Eq) and a near perfect recall of 99.36%. By focusing on the important parts of WebShell behavior, WBFP-Wt makes up for the fact that the smaller model doesn’t understand as much. This results in demonstrations that closely match the desired behavioral profiles, improving both precision and recall.
For GPT-4, WBFP-Wt achieves the highest F1 score of 99.35%, which is 3.03% higher than SC-Sim and 0.34% higher than WBFP-Eq . While precision remains at 100.00% across all WBFP variants, WBFP-Wt increases recall to 98.71%, compared to 92.90% for SC-Sim and 98.03% for WBFP-Eq. This improvement highlights WBFP-Wt’s ability to leverage GPT-4’s advanced understanding of context, matching selected demonstrations to the behavioral characteristics of the target sample to optimize recall without sacrificing precision.
5 Related Work
WebShell Detection Techniques Early efforts in WebShell detection relied on rule-based methods that used signature matching or heuristics to identify malicious scripts (Le et al., 2021; Jinping et al., 2020). For example, Le et al. (2021) proposed H-DLPMWD, a hybrid approach that combines pattern matching with a CNN to detect ASP.NET WebShells, achieving 98.49% accuracy by using Yara-based filtering and opcode indexing. However, such methods struggle against obfuscated or novel WebShell variants due to their reliance on predefined patterns (Hannousse & Yahiouche, 2021). Machine learning (ML) techniques have advanced this landscape by extracting features from code text or runtime behavior. Jinping et al. (2020) introduced a mixed-model approach using Random Forest and CNNs with N-gram and TF-IDF features that achieved 97% accuracy on PHP WebShells, but it requires balanced datasets and struggles with encrypted samples. Deep learning (DL) has further improved adaptability, with models such as CodeBERT for semantic analysis (Pu et al., 2022) and Graph Attention Networks (GAT) for structural insights (Zhang et al., 2025). (Zhang et al., 2025) proposed MMFDetect, which fuses CodeBERT-CL semantics with CNN-extracted visual features from RGB-mapped PHP code, achieving 99.47% accuracy on evasive WebShells. Despite these gains, ML and DL methods require extensive labeled data-a scarcity in cybersecurity-and exhibit limited generalization to obfuscated threats, along with high computational cost (Shang et al., 2024; Jinping et al., 2020).
LLMs in Code Analysis LLMs have revolutionized code-related tasks by leveraging large pre-training corpora for applications such as code generation (Ma et al., 2024), vulnerability detection (Sun et al., 2024), and program reliability assessment (Liu et al., 2024). Ma et al. (2024) demonstrated the ability of LLMs to generate evasive WebShells using hybrid prompts, highlighting their code synthesis potential. Sun et al. (2024) introduced LLM4Vuln, which enhances vulnerability reasoning through knowledge retrieval and has achieved practical success in Solidity audits. These models excel at zero-shot and few-shot learning via prompt engineering (Nong et al., 2024), providing interpretability critical for security contexts (Ma et al., 2024). However, their application to WebShell detection remains underexplored, with previous studies focusing on generation or general vulnerabilities rather than detection of stealthy, context-heavy WebShells.
Challenges of LLMs for WebShell Detection Using LLMs for WebShell detection reveals critical bottlenecks. The fixed context window truncates large WebShells, potentially missing malicious segments embedded in benign code (Ceka et al., 2024; Wang et al., 2025). Fang et al. (2024) found that LLM performance degrades with longer inputs, with GPT-4’s accuracy dropping to 87% on obfuscated JavaScript. Solutions such as chunking or sparse attention (e.g., LongCoder (Guo et al., 2023), SparseCoder (Wang et al., 2024b)) mitigate this, but often lose global context (Wang et al., 2025). In-context learning (ICL), a cornerstone of LLM adaptability, falters because demonstrations consume context space (Min et al., 2022; Wang et al., 2025), and random or semantic similarity-based selections fail to capture WebShell-specific behaviors. Liu & Wang (2023) proposed maximum information gain for ICL, but its focus on text classification limits its applicability to code security.
These gaps between rule-based rigidity, ML and DL data dependency, and LLM context and ICL limitations motivate our BFAD framework. BFAD overcomes context constraints with a hybrid extraction strategy that preserves critical regions and global cues, outperforming naive LLM applications. In addition, our WBFP enhances ICL by prioritizing behaviorally relevant demonstrations, outperforming generic similarity-based selection methods.
Conclusion
This paper presents a novel approach to WebShell detection using LLMs, addressing the unique challenges of applying these models to cybersecurity. We evaluated seven LLMs, including GPT-4 and LLaMA 3.1 70B, against traditional machine learning and deep learning methods using a dataset of 26.59K PHP scripts. Our analysis revealed that larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit high recall but poor precision, both of which underperform state-of-the-art methods such as GAT. To overcome these limitations, we proposed the BFAD framework, which integrates a critical function filter, context-aware code extraction, and WBFP. BFAD significantly improved LLM performance, with GPT-4 achieving an F1 of 99.35%, surpassing traditional benchmarks, and smaller models such as Qwen-2.5-0.5B improving by 51.23 percentage points to an F1 of 82.67%. As the first systematic exploration of LLMs for WebShell detection, this work not only demonstrates their potential, but also provides actionable solutions to their contextual and behavioral challenges, paving the way for future advances in code security.
Appendix A Prompt Details
Prompt for WebShell Detection System Prompt: You are tasked with analyzing PHP scripts. Your objective is to classify the provided PHP code as either a webshell or a legitimate script. A webshell is typically a malicious script intended to exploit the server by executing unauthorized commands or providing backdoor access. User Prompt: Analyze the provided PHP code to determine whether it constitutes a webshell or a legitimate script. Provide your verdict as webshell or benign. [Critical Code] [Source Code] [Examples] Output:
Appendix B Critical Function Details
Table 1: Statistics of Critical Functions in Webshell and Benign Programs. This table reports the percentage of files containing each function category and the average number of occurrences per file, with the “Total” row aggregating statistics across all categories.
Function Category Metric Webshell Files Normal Files
Program Execution Files with Function (%) 53.06 1.54
Avg. Occurrences per File 3.21 0.03
Code Execution Files with Function (%) 85.03 14.79
Avg. Occurrences per File 8.30 0.36
Callback Functions Files with Function (%) 34.69 6.47
Avg. Occurrences per File 0.92 0.11
Network Communication Files with Function (%) 50.34 2.77
Avg. Occurrences per File 1.69 0.04
Information Gathering Files with Function (%) 46.26 2.77
Avg. Occurrences per File 5.46 0.05
Obfuscation and Encryption Files with Function (%) 69.39 9.86
Avg. Occurrences per File 3.19 0.16
Total (All Functions) Files with Function (%) 91.16 20.49
Avg. Occurrences per File 22.76 0.74
Appendix C Dataset Details
Table 2: Dataset Composition, Distribution, and Sources. The dataset comprises 26,594 PHP scripts, categorized into benign programs and webshells, with their respective counts, proportions, and sources.
Category Count Percentage Source References
Benign Programs 21,665 81.5% Grav, OctoberCMS, Laravel, WordPress, Joomla, Nextcloud, Symfony, CodeIgniter, Yii2, CakePHP, Intervention/Image, Typecho
Webshells 4,929 18.5% WebShell, WebshellSample, Awsome-Webshell, PHP-Bypass-Collection, WebShell (tdifg), Webshell (lhlsec), PHP-Backdoors, Tennc/Webshell, PHP-Webshells, BlackArch/Webshells, Webshell-Samples, Programe, WebshellDetection, WebShellCollection, PHP-Backdoors (1337r0j4n), PHP-Webshell-Dataset, Xiao-Webshell
Total 26,594 100.0%
Table 3: Normalized Scores for Key Function Categories. These scores reflect the weighted behavioral significance of each category as computed by the WBFP method.
Function Category Normalized Score
Program Execution 0.2068
Code Execution 0.2081
Callback Functions 0.0790
Network Communication 0.1498
Information Gathering 0.1861
Obfuscation and Encryption 0.1702
Appendix D Results
Table 4: Performance Comparison of BFAD-Enhanced Models Against Baselines. This table compares traditional ML and DL models with large and small-scale LLMs, both standalone and enhanced with BFAD.
Category Model Accuracy Precision Recall F1 Score
Sequence Baselines GloVe+SVM 96.20% 93.30% 94.30% 93.80%
CodeBERT+RF 96.30% 94.00% 95.60% 94.80%
Graph Baselines GCN 96.90% 94.40% 95.30% 94.90%
GAT 98.37% 99.52% 97.39% 98.87%
LLM Baselines (Large) GPT-4 97.27% 100.00% 85.98% 92.46%
LLaMA-3.1-70B 98.01% 97.31% 92.36% 94.77%
Qwen-2.5-Coder-14B 98.64% 99.32% 93.63% 96.39%
LLM Baselines (Small) Qwen-2.5-Coder-3B 71.11% 38.93% 99.32% 55.93%
Qwen-2.5-3B 93.72% 78.03% 91.84% 84.37%
Qwen-2.5-1.5B 43.62% 34.61% 95.77% 50.84%
Qwen-2.5-0.5B 19.47% 18.65% 100.00% 31.44%
LLM + BFAD GPT-4 99.75% 100.00% 98.71% 99.35% (+6.89)
LLaMA-3.1-70B 99.38% 98.72% 98.09% 98.40% (+3.63)
Qwen-2.5-Coder-14B 98.76% 98.68% 94.90% 96.75% (+0.36)
Qwen-2.5-Coder-3B 78.89% 46.67% 100.00% 63.64% (+7.71)
Qwen-2.5-3B 97.39% 88.64% 99.36% 93.69% (+9.32)
Qwen-2.5-1.5B 80.40% 48.51% 100.00% 65.33% (+14.49)
Qwen-2.5-0.5B 91.94% 71.10% 98.73% 82.67% (+51.23)
Table 5: Performance of Context-Aware Code Extraction with Qwen-2.5-3B with Different Context Lengths and Strategies.
Method Accuracy Precision Recall F1 Score
Source Code (Vanilla) 93.72% 78.03% 91.84% 84.37%
Critical Regions ($\tau=100$) 96.28% 86.71% 95.54% 90.91%
Critical Regions ($\tau=200$) 95.78% 84.75% 95.54% 89.82%
Critical Regions ($\tau=300$) 95.04% 82.32% 94.90% 88.17%
Hybrid Strategy ($\tau=100$) 96.40% 89.02% 92.99% 90.97%
Hybrid Strategy ($\tau=200$) 95.66% 85.47% 93.63% 89.36%
Hybrid Strategy ($\tau=300$) 95.78% 85.55% 94.27% 89.70%
Table 6: Performance of Context-Aware Code Extraction with GPT-4 with Different Context Lengths and Strategies.
Method Accuracy Precision Recall F1 Score
Source Code (Vanilla) 97.27% 100.00% 85.98% 92.46%
Critical Regions ($\tau=100$) 98.51% 99.32% 92.99% 96.05%
Critical Regions ($\tau=200$) 99.01% 99.34% 95.54% 97.40%
Critical Regions ($\tau=300$) 99.01% 98.69% 96.18% 97.42%
Hybrid Strategy ($\tau=100$) 99.01% 100.00% 94.90% 97.39%
Hybrid Strategy ($\tau=200$) 99.14% 100.00% 95.81% 97.86%
Hybrid Strategy ($\tau=300$) 99.38% 100.00% 96.82% 98.38%
Table 7: Comparison of Demonstration Selection Strategies for In-Context Learning with Qwen-2.5-3B (Under Best Hybrid Strategy $\tau=100$).
Method Accuracy Precision Recall F1 Score
No-ICL 93.72% 78.03% 91.84% 84.37%
Random 77.79% 46.33% 88.53% 60.83%
SC-Sim 93.42% 78.57% 91.08% 84.36%
WBFP-Eq 96.98% 86.39% 99.32% 92.41%
WBFP-Wt 97.39% 88.64% 99.36% 93.69%
Table 8: Comparison of Demonstration Selection Strategies for In-Context Learning with GPT-4 (Under Best Hybrid Strategy $\tau=300$).
Method Accuracy Precision Recall F1 Score
No-ICL 99.38% 100.00% 96.82% 98.38%
Random 89.00% 65.58% 90.97% 76.22%
SC-Sim 98.63% 100.00% 92.90% 96.32%
WBFP-Eq 99.62% 100.00% 98.03% 99.01%
WBFP-Wt 99.75% 100.00% 98.71% 99.35%

### v2

@@
 Abstract
-WebShell attacks, in which malicious scripts are injected into web servers, are a major cybersecurity threat.
-Traditional machine learning and deep learning methods are hampered by issues such as the need for extensive training data, catastrophic forgetting, and poor generalization.
-Recently, Large Language Models (LLMs) have gained attention for code-related tasks, but their potential in WebShell detection remains underexplored.
+WebShell attacks, where malicious scripts are injected into web servers, pose a significant cybersecurity threat.
+Traditional machine learning and deep learning methods are often hampered by challenges such as the need for extensive training data, catastrophic forgetting, and poor generalization.
+Recently, Large Language Models (LLMs) have emerged as a powerful alternative for code-related tasks, but their potential in WebShell detection remains underexplored.
 In this paper, we make two major contributions: (1) a comprehensive evaluation of seven LLMs, including GPT-4, LLaMA 3.1 70B, and Qwen 2.5 variants, benchmarked against traditional sequence- and graph-based methods using a dataset of 26.59K PHP scripts, and (2) the Behavioral Function-Aware Detection (BFAD) framework, designed to address the specific challenges of applying LLMs to this domain.
 Our framework integrates three components: a Critical Function Filter that isolates malicious PHP function calls, a Context-Aware Code Extraction strategy that captures the most behaviorally indicative code segments, and Weighted Behavioral Function Profiling (WBFP) that enhances in-context learning by prioritizing the most relevant demonstrations based on discriminative function-level profiles.
-Our results show that larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit the opposite trade-off.
-However, all models lag behind previous State-Of-The-Art (SOTA) methods.
-With BFAD, the performance of all LLMs improved, with an average F1 score increase of 13.82%.
-Larger models such as GPT-4, LLaMA 3.1 70B, and Qwen 2.5 14B outperform SOTA methods, while smaller models such as Qwen 2.5 3B achieve performance competitive with traditional methods.
-This work is the first to explore the feasibility and limitations of LLMs for WebShell detection, and provides solutions to address the challenges in this task.
+Our results show that, stemming from their distinct analytical strategies, larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit the opposite trade-off.
+However, all baseline models lag behind previous State-Of-The-Art (SOTA) methods.
+With the application of BFAD, the performance of all LLMs improves significantly, yielding an average F1 score increase of 13.82%.
+Notably, larger models like GPT-4, LLaMA-3.1-70B, and Qwen-2.5-Coder-14B now outperform SOTA benchmarks, while smaller models such as Qwen-2.5-Coder-3B achieve performance competitive with traditional methods.
+This work is the first to explore the feasibility and limitations of LLMs for WebShell detection and provides solutions to address the challenges in this task.
 1 Introduction
-The rapid growth of web applications and the expansion of cloud-based services have escalated the cybersecurity threat landscape, with WebShells emerging as a significant concern.
+The rapid growth of web applications and the expansion of cloud-based services have intensified the cybersecurity threat landscape, with WebShells emerging as a significant concern.
 WebShells are malicious scripts injected into web servers that allow attackers to remotely execute arbitrary commands, steal sensitive data, and compromise system integrity (Starov et al., 2016; Cui et al., 2018).
@@
 Machine learning models, especially deep learning techniques (Pu et al., 2022), have shown promise in addressing these threats.
-However, they require extensive training on large datasets - resources that are often difficult to obtain and sensitive in nature  (Shang et al., 2024).
+However, they require extensive training on large datasets–resources that are often difficult to obtain and sensitive in nature (Shang et al., 2024).
 In addition, these models face challenges such as catastrophic forgetting and poor generalization, especially when dealing with obfuscated or encrypted attacks (Jinping et al., 2020; Zhang et al., 2025).
 Recently, Large Language Models (LLMs) have gained attention for a variety of code-related tasks, including code generation (Ma et al., 2024) and vulnerability detection (Liu & He, 2023; Wang et al., 2025).
-Studies have shown that with timely engineering, LLMs can perform remarkably well without additional training  (Nong et al., 2024; Trad & Chehab, 2025).
+Studies have shown that with effective prompt engineering, LLMs can perform remarkably well without additional training (Nong et al., 2024; Trad & Chehab, 2025).
 They also provide interpretable explanations for their decisions, which is critical in cybersecurity contexts (Ma et al., 2024).
@@
 Detecting WebShells with LLMs presents unique challenges that differ from other code analysis tasks.
-WebShells often employ obfuscation and encryption techniques, and are embedded in large codebases dominated by benign content (Liu & He, 2023).
+WebShells often employ obfuscation and encryption techniques and are embedded in large codebases dominated by benign content (Liu & He, 2023).
 For example, the largest WebShell in our dataset spans 1,386,438 tokens, far exceeding the context window of most LLMs, which risks truncating critical malicious segments when processing entire source files (Wang et al., 2025; Ceka et al., 2024).
-In addition, in-context learning (ICL) struggles in this domain: the variability and obfuscation of WebShells complicates the selection of effective demonstrations, and these examples further occupy significant context space, reducing capacity for the target code (Yuan et al., 2024).
+In addition, in-context learning (ICL) struggles in this domain: the variability and obfuscation of WebShells complicate the selection of effective demonstrations, and these examples further occupy significant context space, reducing capacity for the target code (Yuan et al., 2024).
 While recent research has focused on increasing the context length of LLMs (Chen et al., 2023), studies suggest that the performance of an LLM tends to degrade with longer inputs, and the low processing speed may also become unacceptable for practical use (Ma et al., 2024; Fang et al., 2024).
 In this paper, we present two key contributions to advance the application of LLMs in WebShell detection.
-First, we systematically evaluate LLMs in the context of WebShell detection, comparing their performance with traditional state-of-the-art machine learning (SOTA) methods.
-Specifically, we evaluate seven closed-source and open-source LLMs of different sizes, including GPT-4 (Achiam et al., 2023), LLaMA 3.1 70B (Grattafiori et al., 2024), Qwen 2.5 Coder (14B/3B) (Yang et al., 2024), and Qwen 2.5 (3B/1.5B/0.5B) (Yang et al., 2024) on a dataset containing 26.59K PHP scripts (4.93K WebShells and 21.66K benign samples).
+First, we systematically evaluate LLMs in the context of WebShell detection, comparing their performance with traditional state-of-the-art (SOTA) machine learning methods.
+Specifically, we evaluate seven closed-source and open-source LLMs of different sizes–including GPT-4 (Achiam et al., 2023), LLaMA 3.1 70B (Grattafiori et al., 2024), Qwen 2.5 Coder (14B/3B) (Yang et al., 2024), and Qwen 2.5 (3B/1.5B/0.5B) (Yang et al., 2024)–on a dataset containing 26.59K PHP scripts (4.93K WebShells and 21.66K benign samples).
 Our analysis reveals several key findings:
-  • Larger LLMs, such as GPT-4 and Qwen 2.5 Coder 14B, achieve near perfect precision (close to 100%), but struggle with recall (e.g., GPT-4’s recall is 85.98%), indicating difficulty in detecting all malicious instances.
-  • Smaller LLMs, such as Qwen 2.5 Coder 3B and Qwen 2.5 0.5B, have high recall (close to 100%) but suffer from low precision (e.g.
-Qwen 2.5 Coder 3B’s precision is 38.93%), indicating a tendency to misclassify benign files as malicious.
+  • Larger LLMs, such as GPT-4 and Qwen 2.5 Coder 14B, achieve near-perfect precision but struggle with recall (e.g., GPT-4’s recall is 85.98%).
+Their deep contextual analysis, while powerful, can be deceived by well-disguised malicious code, leading to false negatives.
+  • Smaller LLMs, such as Qwen 2.5 Coder 3B and Qwen 2.5 0.5B, exhibit high recall but suffer from low precision (e.g., Qwen 2.5 Coder 3B’s precision is 38.93%).
+This stems from their reliance on surface-level pattern matching, which often flags any use of sensitive functions without sufficient contextual validation.
   • Randomly selected ICL demonstrations degrade LLM detection performance; examples selected based on semantic similarity to the source code do not yield significant improvements.
-We compare the performance of LLMs with traditional methods, including Glove+SVM (Petridis, 2024; Rigutini et al., 2024), CodeBERT+Random Forest (Alghamdi et al., 2022), and graph-based approaches such as GAT (Kang et al., 2023).
-The best performing LLM, Qwen 2.5 Coder 14B, achieves an F1 score of 96.39%, although it still lags behind GAT-based methods, which achieve an F1 score of 98.87%.
+For this comparison, we benchmark against several traditional methods, including GloVe+SVM (Petridis, 2024; Rigutini et al., 2024), CodeBERT+Random Forest (Alghamdi et al., 2022), and graph-based approaches such as GAT (Kang et al., 2023).
+The best-performing LLM, Qwen 2.5 Coder 14B, achieves an F1 score of 96.39%, although it still lags behind GAT-based methods, which achieve an F1 score of 98.87%.
 [^1] While traditional models such as GAT have demonstrated superior performance, they require extensive training and significant data collection efforts.
-In contrast, LLMs enable direct detection through prompts, leveraging their pre-trained knowledge with minimal task-specific resources to deliver competitive results..
-Second, we present the Behavioral Function-Aware Detection (BFAD) framework to address the identified challenges and to improve the performance of LLMs to meet the requirements of downstream applications by achieving a more balanced trade-off between precision and recall.
+In contrast, LLMs enable direct detection through prompts, leveraging their pre-trained knowledge with minimal task-specific resources to deliver competitive results.
+Second, we present the Behavioral Function-Aware Detection (BFAD) framework to address the identified challenges and improve the performance of LLMs to meet the requirements of downstream applications by achieving a more balanced trade-off between precision and recall.
 Our framework combines risk-filtering techniques with an enhanced ICL strategy that uses weighted demonstration selection to prioritize examples most closely related to key malicious behaviors.
-Experimental results show that our approach improves the average F1 score across all LLMs by 13.82%, with GPT-4 and Qwen 2.5 0.5B improving by 6.89% and 51.32%, respectively.
-In several cases, such as GPT-4, LLaMA 3.1 70B, Qwen 2.5 Coder 14B, and Qwen 2.5 Coder 3B, our approach enables performance that is competitive or even superior to traditional methods.
+Experimental results show that our approach improves the average F1 score across all LLMs by 13.82%, with GPT-4 and Qwen 2.5 0.5B improving by 6.89% and 51.23%, respectively.
+In several cases, such as GPT-4, LLaMA 3.1 70B, Qwen 2.5 Coder 14B, and Qwen 2.5 Coder 3B, our approach enables performance that is competitive with or even superior to traditional methods.
 To the best of our knowledge, this is the first work to systematically analyze the feasibility and limitations of applying LLMs to WebShell detection.
@@
 6:  for each occurrence of $f$ at position $p$ do
-7:        Extract context window $[p-\tau,p+\tau]$ from $\mathcal{C}$
+7:   Extract context window $p-\tau,p+\tau]$ from $\mathcal{C}$
 8:   Add the extracted region to $\mathcal{R}$
@@
 ICL Settings.
-We randomly selected 60% of the dataset to create a fixed known demonstration library for ICL.
+We randomly selected 60% of the dataset to create a fixed, known demonstration library for ICL.
 Using this subset, we computed normalized scores for different function categories based on the WBFP method, giving equal weight to coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$) to profile functions according to their behavioral importance in distinguishing WebShells from benign programs.
@@
 4 Results and Analysis
-In this section, we systematically evaluate the performance of LLMs of different scales for WebShell detection, and assess the improvements provided by our proposed BFAD framework.
+In this section, we systematically evaluate the performance of LLMs of different scales for WebShell detection and assess the improvements provided by our proposed BFAD framework.
 Our analysis addresses three research questions (RQs) to explore both the baseline LLM capabilities and the effectiveness of the BFAD components:
-  • RQ1: How do large and small scale LLMs perform in WebShell detection compared to traditional ML and DL methods, and how does BFAD improve their effectiveness?
+  • RQ1: How do large and small-scale LLMs perform in WebShell detection compared to traditional ML and DL methods, and how does BFAD improve their effectiveness?
   • RQ2: How effective is context-aware code extraction at balancing global context and local behavioral focus under LLM context length constraints?
@@
 4.1 Performance Evaluation of LLMs and BFAD Enhancements (RQ1)
-We evaluated seven LLMs, including large-scale models such as GPT-4, LLaMA-3.1-70B, and Qwen-2.5-Coder-14B, and small-scale models such as Qwen-2.5-Coder-3B, Qwen-2.5-3B, Qwen-2.5-1.5B, and Qwen-2.5-0.5B.
-These were compared with traditional ML and DL baselines, including sequence-based methods (Glove+SVM, CodeBERT+RF) and graph-based methods (GCN, GAT).
-The results are detailed in Table 4 in Appendix D.
-Characteristics of LLMs in WebShell Detection
-Our evaluation reveals distinct performance characteristics of LLMs in WebShell detection that are influenced by their size.
-Compared to traditional machine learning and deep learning methods, especially GAT, vanilla LLMs exhibit unbalanced performance - large models lack sufficient recall, while small models lack precision - making them less effective without targeted improvements.
-In particular, large LLMs such as GPT-4 and Qwen-2.5-Coder-14B achieve high precision (100% and 99.32%, respectively) but moderate recall (85.98% and 93.63%, respectively), reflecting a conservative classification bias.
-This is likely due to their extensive training on diverse datasets, which prioritizes accurate identification of benign code over detection of rare or novel WebShells.
-Code-specific fine-tuning improves the performance of Qwen-2.5-Coder-14B over general-purpose models such as LLaMA-3.1-70B (97.31% precision, 92.36% recall), underscoring the benefits of domain specialization.
-In contrast, small-scale LLMs such as Qwen-2.5-0.5B and Qwen-2.5-1.5B exhibit nearly perfect recall (100% and 95.77%, respectively) but extremely low precision (18.65% and 34.61%, respectively).
-This imbalance may be due to their increased sensitivity to prompts, making them more likely to provide answers that match user expectations, a problem exacerbated by their limited ability to model complex code relationships.
-Effects of BFAD on LLMs
-The BFAD framework significantly improves the performance of LLMs for WebShell detection.
-For large-scale LLMs, BFAD increases recall while maintaining near-perfect precision; in particular, GPT-4’s recall increases by 12.73% to 98.71%, yielding an F1 score of 99.35%, which exceeds the graph-based GAT baseline (98.87% F1), while LLaMA-3.1-70B achieves an F1 score of 98.40%, closely matching GAT.
-Conversely, for small LLMs, BFAD significantly increases precision without compromising high recall; Qwen-2.5-0.5B’s precision increases by 52.36% to 71.10%, resulting in an F1 score of 82.67% (a 51.23% improvement), and Qwen-2.5-1.5B achieves an F1 score of 65.33% (up 14.85%).
-These advances enable BFAD-enhanced LLMs to match or exceed traditional methods without the need for additional fine-tuning or training, driven by a strategic focus on critical code regions and the use of weighted behavioral profiling to select contextually relevant demonstrations.
-This approach effectively mitigates the inherent context length constraints of LLMs and improves their generalization from limited examples.
-Notably, small-scale LLMs exhibit more pronounced gains, underscoring the critical role of structured input for models with limited capacity, while large-scale LLMs benefit from improved recall, making them suitable for safety-critical applications.
-These results highlight the transformative potential of integrating domain-specific strategies with LLMs to address specialized challenges, such as WebShell detection, with exceptional effectiveness.
+To answer our first research question, we evaluated seven LLMs against traditional ML and DL baselines.
+The LLMs included large-scale models (GPT-4, LLaMA-3.1-70B, Qwen-2.5-Coder-14B) and small-scale models (Qwen-2.5-Coder-3B, Qwen-2.5-3B, Qwen-2.5-1.5B, Qwen-2.5-0.5B).
+Baselines consisted of sequence-based methods (Glove+SVM, CodeBERT+RF) and graph-based methods (GCN, GAT).
+The detailed performance comparison is presented in Table 1.
+Table 1: Performance Comparison of BFAD-Enhanced Models Against Baselines.
+Category Model Accuracy Precision Recall F1 Score
+Sequence Baselines GloVe+SVM 96.20% 93.30% 94.30% 93.80%
+CodeBERT+RF 96.30% 94.00% 95.60% 94.80%
+Graph Baselines GCN 96.90% 94.40% 95.30% 94.90%
+GAT 98.37% 99.52% 97.39% 98.87%
+LLM Baselines (Large) GPT-4 97.27% 100.00% 85.98% 92.46%
+LLaMA-3.1-70B 98.01% 97.31% 92.36% 94.77%
+Qwen-2.5-Coder-14B 98.64% 99.32% 93.63% 96.39%
+LLM Baselines (Small) Qwen-2.5-Coder-3B 71.11% 38.93% 99.32% 55.93%
+Qwen-2.5-3B 93.72% 78.03% 91.84% 84.37%
+Qwen-2.5-1.5B 43.62% 34.61% 95.77% 50.84%
+Qwen-2.5-0.5B 19.47% 18.65% 100.00% 31.44%
+LLM + BFAD GPT-4 99.75% 100.00% 98.71% 99.35% (+6.89)
+LLaMA-3.1-70B 99.38% 98.72% 98.09% 98.40% (+3.63)
+Qwen-2.5-Coder-14B 98.76% 98.68% 94.90% 96.75% (+0.36)
+Qwen-2.5-Coder-3B 78.89% 46.67% 100.00% 63.64% (+7.71)
+Qwen-2.5-3B 97.39% 88.64% 99.36% 93.69% (+9.32)
+Qwen-2.5-1.5B 80.40% 48.51% 100.00% 65.33% (+14.49)
+Qwen-2.5-0.5B 91.94% 71.10% 98.73% 82.67% (+51.23)
+Performance Summary
+Our experimental results yield two core findings.
+First, vanilla LLMs exhibit a stark, size-dependent trade-off between precision and recall, with all models lagging behind the state-of-the-art GAT baseline.
+For instance, GPT-4 achieves 100% precision but only 85.98% recall, while the small Qwen-2.5-0.5B model reaches 100% recall but a mere 18.65% precision.
+This indicates that out-of-the-box LLMs are not well-calibrated for the nuanced task of WebShell detection, making targeted intervention necessary for effective deployment.
+Second, the BFAD framework dramatically improves performance across all models, elevating large LLMs to surpass the GAT baseline and making small models viable competitors.
+With BFAD, GPT-4’s F1 score increases by 6.89 points to 99.35%, exceeding the GAT baseline.
+The most significant gain is seen in Qwen-2.5-0.5B, whose F1 score improves by a remarkable 51.23 points.
+This demonstrates that by strategically guiding the model’s focus on behaviorally relevant code, BFAD effectively bridges the performance gap, unlocking the potential of LLMs for this task without resource-intensive retraining.
+Analysis of Model Behavior: Misclassification Patterns
+The performance disparities observed in our results appear to stem from fundamental differences in the models’ analytical reasoning.
+The precision-recall imbalance in baseline LLMs is a direct reflection of their reasoning strategy.
+Small models like Qwen-2.5-0.5B tend to rely on surface-level heuristics; they identify keywords like base64_decode and flag them as malicious without assessing the surrounding code’s intent.
+This leads to high recall but generates numerous false positives when such functions are used legitimately (e.g., in a cryptography library).
+In contrast, large models like LLaMA-3.1-70B employ deep contextual reasoning, correctly identifying legitimate use cases and thus achieving high precision.
+However, this sophistication can be a double-edged sword, as these models can be misled by well-disguised WebShells, rationalizing away malicious indicators and causing false negatives that reduce recall.
+Our proposed BFAD mitigates these issues by guiding smaller models to better understand function-level behaviors and helping larger models focus on critical functions.
+As a result, it significantly improves detection performance across scales.
+Analysis of Model Behavior: Counterintuitive Scaling
+We observed a counterintuitive scaling effect where the Qwen-2.5-1.5B model, despite outperforming the 0.5B variant in baseline tests, performed worse with BFAD.
+We attribute this to differing sensitivities to ICL across scales.
+The smallest models, like the 0.5B variant, are highly susceptible to label bias and tend to mimic the ICL demonstration’s label without deep reasoning (Fei et al., 2023).
+Because BFAD provides accurately matched demonstrations via WBFP, the 0.5B model benefits disproportionately from this behavior.
+In contrast, the 1.5B model begins to exhibit more independent reasoning based on its pre-training knowledge.
+This emergent but still imperfect reasoning can conflict with the ICL example’s guidance, degrading performance.
+This aligns with prior findings that LLMs sometimes rely more on pre-training than on provided examples (Peng et al., 2025).
+Therefore, the 0.5B model’s superior performance with BFAD does not indicate greater intelligence but is an artifact of its strong alignment with the high-quality examples provided.
 4.2 The Effectiveness of Context-Aware Code Extraction (RQ2)
@@
 This evaluation builds on the Context-Aware Code Extraction with the hybrid strategy.
-Five demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt).
+Four demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt).
 Results are detailed in Tables 7 and 8 in Appendix D.
@@
 This is mainly because there is a large drop in precision from 78.03% to 46.33%, although recall remains strong at 88.53%.
-Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 22.16% decrease from the No-ICL baseline of 98.38%, driven by a drop in precision from 100.00% to 65.58%.
+Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 16.24% decrease from the No-ICL baseline of 92.46%, driven by a drop in precision from 100.00% to 65.58%.
 These results indicate that Random Selection does not provide contextually relevant demonstrations, making it ineffective for WebShell detection.
@@
 For Qwen-2.5-3B, SC-Sim achieves an F1 score of 84.36%, which is almost the same as the No-ICL baseline (84.37%), with a precision of 78.57% and a recall of 91.08%.
-For GPT-4, it achieves an F1 score of 96.32%, with perfect precision (100.00%), but a reduced recall of 92.90% compared to the No-ICL baseline’s 96.82%.
+For GPT-4, it achieves an F1 score of 96.32%, with perfect precision, but a reduced recall of 92.90% compared to the No-ICL baseline’s 85.98%.
 This limited performance is likely due to the dominance of behaviorally irrelevant code segments in the similarity calculation, which dilutes the focus on critical behavioral patterns essential for accurate WebShell identification.
@@
 In addition, our WBFP enhances ICL by prioritizing behaviorally relevant demonstrations, outperforming generic similarity-based selection methods.
-Conclusion
+6 Limitations and Future Work
+Our research highlights several limitations that present important directions for future investigation in this emerging field.
+6.1 Dataset Generalization
+A fundamental challenge in WebShell detection is the scarcity of real-world data due to privacy concerns.
+Current research, including this study, mainly relies on public datasets, which may not fully represent the diversity and sophistication of novel threats.
+This reliance makes it difficult to rigorously evaluate a model’s ability to generalize to unseen attack vectors.
+This limits rigorous evaluation of a model’s ability to generalize against novel and heavily obfuscated threats.
+To address this, future work should focus on using LLMs to generate synthetic benchmarks.
+A large-scale, LLM-generated dataset would provide a privacy-preserving and diverse alternative for stress-testing model robustness against a wider range of attack vectors.
+6.2 Model Robustness and System Adaptability
+While our BFAD framework is effective, the underlying static analysis approach is vulnerable to sophisticated obfuscation and zero-day threats.
+Enhancing the core model and building adaptive deployment systems are critical next steps.
+To improve model robustness, future work could integrate dynamic analysis, examining runtime function call traces from sandboxed execution to unmask behavior hidden by obfuscation.
+Additionally, since our analysis revealed semantic understanding bottlenecks even in large models, domain-specific fine-tuning of code-centric LLMs on curated WebShell datasets is a crucial step to deepen their expertise and improve detection accuracy.
+For practical and resilient deployment, these enhanced models could be integrated into advanced architectures.
+A multi-stage ”fast and slow” system could use a smaller LLM for rapid, efficient filtering, while reserving a larger reasoning model for in-depth analysis of high-risk samples.
+To ensure long-term effectiveness against evolving threats, this pipeline could be governed by an autonomous agent that continuously identifies new attack patterns and automatically updates the detection system.
+7 Conclusion
 This paper presents a novel approach to WebShell detection using LLMs, addressing the unique challenges of applying these models to cybersecurity.
@@
 Appendix D Results
-Table 4: Performance Comparison of BFAD-Enhanced Models Against Baselines. This table compares traditional ML and DL models with large and small-scale LLMs, both standalone and enhanced with BFAD.
-Category Model Accuracy Precision Recall F1 Score
-Sequence Baselines GloVe+SVM 96.20% 93.30% 94.30% 93.80%
-CodeBERT+RF 96.30% 94.00% 95.60% 94.80%
-Graph Baselines GCN 96.90% 94.40% 95.30% 94.90%
-GAT 98.37% 99.52% 97.39% 98.87%
-LLM Baselines (Large) GPT-4 97.27% 100.00% 85.98% 92.46%
-LLaMA-3.1-70B 98.01% 97.31% 92.36% 94.77%
-Qwen-2.5-Coder-14B 98.64% 99.32% 93.63% 96.39%
-LLM Baselines (Small) Qwen-2.5-Coder-3B 71.11% 38.93% 99.32% 55.93%
-Qwen-2.5-3B 93.72% 78.03% 91.84% 84.37%
-Qwen-2.5-1.5B 43.62% 34.61% 95.77% 50.84%
-Qwen-2.5-0.5B 19.47% 18.65% 100.00% 31.44%
-LLM + BFAD GPT-4 99.75% 100.00% 98.71% 99.35% (+6.89)
-LLaMA-3.1-70B 99.38% 98.72% 98.09% 98.40% (+3.63)
-Qwen-2.5-Coder-14B 98.76% 98.68% 94.90% 96.75% (+0.36)
-Qwen-2.5-Coder-3B 78.89% 46.67% 100.00% 63.64% (+7.71)
-Qwen-2.5-3B 97.39% 88.64% 99.36% 93.69% (+9.32)
-Qwen-2.5-1.5B 80.40% 48.51% 100.00% 65.33% (+14.49)
-Qwen-2.5-0.5B 91.94% 71.10% 98.73% 82.67% (+51.23)
 Table 5: Performance of Context-Aware Code Extraction with Qwen-2.5-3B with Different Context Lengths and Strategies.
@@
 Method Accuracy Precision Recall F1 Score
-No-ICL 99.38% 100.00% 96.82% 98.38%
+No-ICL 97.27% 100.00% 85.98% 92.46%
 Random 89.00% 65.58% 90.97% 76.22%

### v3

@@
 At $\tau=100$, the F1 score increases from 84.37% to 90.91% (+6.54%), with precision increasing from 78.03% to 86.71% (+8.68%) and recall increasing from 91.84% to 95.54% (+3.70%).
-This improvement validates the focus on behaviorally relevant code snippets, which reduces irrelevant context and sharpen the focus of the model.
+This improvement validates the focus on behaviorally relevant code snippets, which reduces irrelevant context and sharpens the focus of the model.
 However, as the context length ($\tau$) increases, performance decreases - F1 drops to 88.17% at $\tau=300$, likely due to the model’s limited ability to handle extended context, which introduces noise that degrades accuracy.
@@
 We evaluated the effectiveness of WBFP for ICL demonstration selection using Qwen-2.5-3B and GPT-4.
-This evaluation builds on the Context-Aware Code Extraction with the hybrid strategy.
+This evaluation builds on the optimal hybrid strategy from RQ2.
 Four demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt).
@@
 Random selection makes ICL performance much worse by adding irrelevant examples, which makes both models much less effective.
-For Qwen-2.5-3B, the F1 score drops to 60.83% under the hybrid strategy with a context length of $\tau=100$, which is a 23.54% reduction from the no-ICL baseline of 84.37%.
-This is mainly because there is a large drop in precision from 78.03% to 46.33%, although recall remains strong at 88.53%.
-Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 16.24% decrease from the No-ICL baseline of 92.46%, driven by a drop in precision from 100.00% to 65.58%.
+For Qwen-2.5-3B, the F1 score drops to 60.83% under the hybrid strategy with a context length of $\tau=100$, which is a 30.14% reduction from the no-ICL baseline of 90.97%.
+This is mainly because there is a large drop in precision from 89.02% to 46.33%, although recall remains strong at 88.53%.
+Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 22.16% decrease from the No-ICL baseline of 98.38%, driven by a drop in precision from 100.00% to 65.58%.
 These results indicate that Random Selection does not provide contextually relevant demonstrations, making it ineffective for WebShell detection.
 The SC-Sim approach, which relies on semantic similarity calculated over entire source code samples, also doesn’t work well.
-For Qwen-2.5-3B, SC-Sim achieves an F1 score of 84.36%, which is almost the same as the No-ICL baseline (84.37%), with a precision of 78.57% and a recall of 91.08%.
-For GPT-4, it achieves an F1 score of 96.32%, with perfect precision, but a reduced recall of 92.90% compared to the No-ICL baseline’s 85.98%.
+For Qwen-2.5-3B, SC-Sim achieves an F1 score of 84.36%, a notable decrease from the No-ICL baseline of 90.97%, with a precision of 78.57% and a recall of 91.08%.
+For GPT-4, it achieves an F1 score of 96.32%, with perfect precision, but a reduced recall of 92.90% compared to the No-ICL baseline’s 96.82%.
 This limited performance is likely due to the dominance of behaviorally irrelevant code segments in the similarity calculation, which dilutes the focus on critical behavioral patterns essential for accurate WebShell identification.
@@
 Specifically, for Qwen-2.5-3B, WBFP-Wt achieves an F1 score of 93.69%, which is 9.33% higher than SC-Sim and 1.28% higher than WBFP-Eq.
-This is because WBFP-Wt achieves a precision of 88.64% (compared to 78.57% for SC-Sim and 86.39% for WBFP-Eq) and a near perfect recall of 99.36%.
+This is because WBFP-Wt achieves a precision of 88.64% (compared to 78.57% for SC-Sim and 86.39% for WBFP-Eq) and a near-perfect recall of 99.36%.
 By focusing on the important parts of WebShell behavior, WBFP-Wt makes up for the fact that the smaller model doesn’t understand as much.
@@
 This paper presents a novel approach to WebShell detection using LLMs, addressing the unique challenges of applying these models to cybersecurity.
-We evaluated seven LLMs, including GPT-4 and LLaMA 3.1 70B, against traditional machine learning and deep learning methods using a dataset of 26.59K PHP scripts.
+We evaluated seven LLMs against traditional machine learning and deep learning methods using a dataset of 26.59K PHP scripts.
 Our analysis revealed that larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit high recall but poor precision, both of which underperform state-of-the-art methods such as GAT.
@@
 BFAD significantly improved LLM performance, with GPT-4 achieving an F1 of 99.35%, surpassing traditional benchmarks, and smaller models such as Qwen-2.5-0.5B improving by 51.23 percentage points to an F1 of 82.67%.
-As the first systematic exploration of LLMs for WebShell detection, this work not only demonstrates their potential, but also provides actionable solutions to their contextual and behavioral challenges, paving the way for future advances in code security.
+As the first systematic exploration of LLMs for WebShell detection, this work provides actionable solutions to their contextual and behavioral challenges, paving the way for future advances in code security.
 Appendix A Prompt Details
@@
 Method Accuracy Precision Recall F1 Score
-No-ICL 93.72% 78.03% 91.84% 84.37%
+No-ICL 96.40% 89.02% 92.99% 90.97%
 Random 77.79% 46.33% 88.53% 60.83%
@@
 Method Accuracy Precision Recall F1 Score
-No-ICL 97.27% 100.00% 85.98% 92.46%
+No-ICL 99.38% 100.00% 96.82% 98.38%
 Random 89.00% 65.58% 90.97% 76.22%

### v4

@@
 Finally, Obfuscation and Encryption functions, such as base64_encode and openssl_encrypt, help disguise or encrypt malicious payloads to avoid detection.
-Our statistical analysis (detailed in Appendix B) shows that WebShell files use critical functions far more often than benign files.
+Our statistical analysis (detailed in Appendix C) shows that WebShell files use critical functions far more often than benign files.
 On average, WebShells contain 22.76 calls to critical functions, compared to only 0.74 in benign files, underscoring their behavioral complexity and distinctiveness.
@@
 To balance efficiency and performance, we limit the user query to one ICL demonstration, which reduces computational overhead while preserving sufficient context for reliable detection.
-Our prompt is described in detail in Appendix A.
+Our prompt is described in detail in Appendix B.
 3 Experiment
@@
 These results indicate that WebShells are typically significantly longer than benign samples.
-A detailed summary of the dataset composition can be found in the Table 3 in the Appendix C.
+A detailed summary of the dataset composition can be found in the Table 3 in the Appendix D.
 3.2 Experiment Setup
@@
 Using this subset, we computed normalized scores for different function categories based on the WBFP method, giving equal weight to coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$) to profile functions according to their behavioral importance in distinguishing WebShells from benign programs.
-These scores, detailed in the Table 4 in the Appendix C, guided the selection of ICL demonstrations from the known sample library.
+These scores, detailed in the Table 4 in the Appendix D, guided the selection of ICL demonstrations from the known sample library.
 Baseline Models, Hyperparameter Settings, and Evaluation Metrics.
@@
 Three configurations were compared: (1) predictions based on the full source code, (2) predictions using only extracted critical regions, and (3) a hybrid approach combining critical regions with truncated source code.
-Results are reported in Tables 5 and 6 in Appendix D.
+Results are reported in Tables 5 and 6 in Appendix E.
 Impact of Critical Regions
@@
 Four demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt).
-Results are detailed in Tables 7 and 8 in Appendix D.
+Results are detailed in Tables 7 and 8 in Appendix E.
 Limitations of Random and Semantic Similarity-Based Selection
@@
 In addition, our WBFP enhances ICL by prioritizing behaviorally relevant demonstrations, outperforming generic similarity-based selection methods.
-6 Limitations and Future Work
-Our research highlights several limitations that present important directions for future investigation in this emerging field.
-6.1 Dataset Generalization
-A fundamental challenge in WebShell detection is the scarcity of real-world data due to privacy concerns.
-Current research, including this study, mainly relies on public datasets, which may not fully represent the diversity and sophistication of novel threats.
-This reliance makes it difficult to rigorously evaluate a model’s ability to generalize to unseen attack vectors.
-This limits rigorous evaluation of a model’s ability to generalize against novel and heavily obfuscated threats.
-To address this, future work should focus on using LLMs to generate synthetic benchmarks.
+6 Conclusion
+This paper systematically demonstrates that baseline LLMs are suboptimal for WebShell detection, exhibiting a precision-recall trade-off that underperforms state-of-the-art methods.
+We propose the Behavioral Function-Aware Detection (BFAD) framework, which overcomes these limitations by focusing analysis on critical code segments and optimizing in-context learning.
+BFAD enables our LLM-based approach to surpass traditional benchmarks.
+Appendix A Limitations and Future Work
+Our research also highlights several limitations that present important directions for future investigation in this emerging field.
+A.1 Dataset Generalization
+A fundamental challenge in WebShell research is the scarcity of diverse, real-world data.
+Our reliance on public datasets, like that of prior work, limits the evaluation of model generalization against novel, sophisticated threats.
+To address this, future work should focus on using LLMs to generate large-scale synthetic benchmarks.
 A large-scale, LLM-generated dataset would provide a privacy-preserving and diverse alternative for stress-testing model robustness against a wider range of attack vectors.
-6.2 Model Robustness and System Adaptability
-While our BFAD framework is effective, the underlying static analysis approach is vulnerable to sophisticated obfuscation and zero-day threats.
-Enhancing the core model and building adaptive deployment systems are critical next steps.
-To improve model robustness, future work could integrate dynamic analysis, examining runtime function call traces from sandboxed execution to unmask behavior hidden by obfuscation.
-Additionally, since our analysis revealed semantic understanding bottlenecks even in large models, domain-specific fine-tuning of code-centric LLMs on curated WebShell datasets is a crucial step to deepen their expertise and improve detection accuracy.
-For practical and resilient deployment, these enhanced models could be integrated into advanced architectures.
-A multi-stage ”fast and slow” system could use a smaller LLM for rapid, efficient filtering, while reserving a larger reasoning model for in-depth analysis of high-risk samples.
-To ensure long-term effectiveness against evolving threats, this pipeline could be governed by an autonomous agent that continuously identifies new attack patterns and automatically updates the detection system.
-7 Conclusion
-This paper presents a novel approach to WebShell detection using LLMs, addressing the unique challenges of applying these models to cybersecurity.
-We evaluated seven LLMs against traditional machine learning and deep learning methods using a dataset of 26.59K PHP scripts.
-Our analysis revealed that larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit high recall but poor precision, both of which underperform state-of-the-art methods such as GAT.
-To overcome these limitations, we proposed the BFAD framework, which integrates a critical function filter, context-aware code extraction, and WBFP.
-BFAD significantly improved LLM performance, with GPT-4 achieving an F1 of 99.35%, surpassing traditional benchmarks, and smaller models such as Qwen-2.5-0.5B improving by 51.23 percentage points to an F1 of 82.67%.
-As the first systematic exploration of LLMs for WebShell detection, this work provides actionable solutions to their contextual and behavioral challenges, paving the way for future advances in code security.
-Appendix A Prompt Details
+A.2 Model and System Robustness
+Our framework’s static analysis approach remains vulnerable to advanced obfuscation and zero-day attacks.
+Future work should therefore focus on enhancing model robustness and system adaptability.
+To counter evasion, models could integrate dynamic analysis by examining runtime function call traces from sandboxed execution.
+To deepen semantic understanding, domain-specific fine-tuning of code-centric LLMs on curated WebShell datasets is a crucial next step.
+For practical deployment, these enhanced models could be integrated into a multi-stage ”fast and slow” system, using a smaller model for rapid filtering and a larger one for in-depth analysis of high-risk samples.
+To ensure long-term resilience, this pipeline could be governed by an autonomous agent that continuously identifies new attack patterns and automatically updates the detection models.
+Appendix B Prompt Details
 Prompt for WebShell Detection System Prompt: You are tasked with analyzing PHP scripts.
@@
 [Critical Code] [Source Code] [Examples] Output:
-Appendix B Critical Function Details
+Appendix C Critical Function Details
 Table 2: Statistics of Critical Functions in WebShell and Benign Programs.
@@
 Avg. Occurrences per File 22.76 0.74
-Appendix C Dataset Details
+Appendix D Dataset Details
 Table 3: Dataset Composition, Distribution, and Sources.
@@
 Obfuscation and Encryption 0.1702
-Appendix D Results
+Appendix E Results
 Table 5: Performance of Context-Aware Code Extraction with Qwen-2.5-3B with Different Context Lengths and Strategies.

### v5

@@
 Abstract
-WebShell attacks, where malicious scripts are injected into web servers, pose a significant cybersecurity threat.
-Traditional machine learning and deep learning methods are often hampered by challenges such as the need for extensive training data, catastrophic forgetting, and poor generalization.
-Recently, Large Language Models (LLMs) have emerged as a powerful alternative for code-related tasks, but their potential in WebShell detection remains underexplored.
-In this paper, we make two major contributions: (1) a comprehensive evaluation of seven LLMs, including GPT-4, LLaMA 3.1 70B, and Qwen 2.5 variants, benchmarked against traditional sequence- and graph-based methods using a dataset of 26.59K PHP scripts, and (2) the Behavioral Function-Aware Detection (BFAD) framework, designed to address the specific challenges of applying LLMs to this domain.
-Our framework integrates three components: a Critical Function Filter that isolates malicious PHP function calls, a Context-Aware Code Extraction strategy that captures the most behaviorally indicative code segments, and Weighted Behavioral Function Profiling (WBFP) that enhances in-context learning by prioritizing the most relevant demonstrations based on discriminative function-level profiles.
-Our results show that, stemming from their distinct analytical strategies, larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit the opposite trade-off.
-However, all baseline models lag behind previous State-Of-The-Art (SOTA) methods.
-With the application of BFAD, the performance of all LLMs improves significantly, yielding an average F1 score increase of 13.82%.
-Notably, larger models like GPT-4, LLaMA-3.1-70B, and Qwen-2.5-Coder-14B now outperform SOTA benchmarks, while smaller models such as Qwen-2.5-Coder-3B achieve performance competitive with traditional methods.
-This work is the first to explore the feasibility and limitations of LLMs for WebShell detection and provides solutions to address the challenges in this task.
+WebShell attacks—where adversaries implant malicious scripts on web servers—remain a persistent threat.
+Prior machine-learning and deep-learning detectors typically depend on task-specific supervision and can be brittle under data scarcity, rapid concept drift, and out-of-distribution (OOD) deployment.
+Large language models (LLMs) have recently shown strong code understanding capabilities, but their reliability for WebShell detection remains unclear.
+We address this gap by (i) systematically evaluating seven LLMs (including GPT-4, LLaMA-3.1-70B, and Qwen-2.5 variants) against representative sequence- and graph-based baselines on 26.59K PHP scripts, and (ii) proposing Behavioral Function-Aware Detection (BFAD), a behavior-centric framework that adapts LLM inference to WebShell-specific execution patterns.
+BFAD anchors analysis on security-sensitive PHP functions via a Critical Function Filter, constructs compact LLM inputs with Context-Aware Code Extraction, and selects in-context demonstrations using Weighted Behavioral Function Profiling (WBFP), which ranks examples by a behavior-weighted, function-level similarity.
+Empirically, we observe a consistent precision–recall asymmetry: larger LLMs often achieve high precision but miss attacks (lower recall), while smaller models exhibit the opposite tendency; moreover, off-the-shelf LLM prompting underperforms established detectors.
+BFAD substantially improves all evaluated LLMs, boosting F1 by 13.82% on average; notably, GPT-4, LLaMA-3.1-70B, and Qwen-2.5-Coder-14B exceed prior SOTA benchmarks, while Qwen-2.5-Coder-3B becomes competitive with traditional methods.
+Overall, our results clarify when LLMs succeed or fail on WebShell detection, provide a practical recipe and highlight future directions for making LLM-based detection more reliable.
 1 Introduction
-The rapid growth of web applications and the expansion of cloud-based services have intensified the cybersecurity threat landscape, with WebShells emerging as a significant concern.
-WebShells are malicious scripts injected into web servers that allow attackers to remotely execute arbitrary commands, steal sensitive data, and compromise system integrity (Starov et al., 2016; Cui et al., 2018).
-According to a Cisco Talos report (Talos, 2024), threat groups used a variety of web shells against vulnerable or unpatched web applications in 35% of incidents in Q4 2024, a sharp increase from the previous quarter, when such activity was observed in only 10% of incidents.
-These attacks are particularly dangerous due to the stealthy nature of WebShells, which are rapidly evolving to evade traditional detection methods (Hannousse & Yahiouche, 2021).
-In response to this challenge, the community has proposed several approaches.
-Rule-based methods that rely on predefined signatures or heuristics are increasingly ineffective against the complexity and diversity of modern WebShells (Le et al., 2021; Jinping et al., 2020).
-Machine learning models, especially deep learning techniques (Pu et al., 2022), have shown promise in addressing these threats.
-However, they require extensive training on large datasets–resources that are often difficult to obtain and sensitive in nature (Shang et al., 2024).
-In addition, these models face challenges such as catastrophic forgetting and poor generalization, especially when dealing with obfuscated or encrypted attacks (Jinping et al., 2020; Zhang et al., 2025).
-Recently, Large Language Models (LLMs) have gained attention for a variety of code-related tasks, including code generation (Ma et al., 2024) and vulnerability detection (Liu & He, 2023; Wang et al., 2025).
-Studies have shown that with effective prompt engineering, LLMs can perform remarkably well without additional training (Nong et al., 2024; Trad & Chehab, 2025).
-They also provide interpretable explanations for their decisions, which is critical in cybersecurity contexts (Ma et al., 2024).
-Despite these advantages, there has been limited exploration of LLMs for WebShell detection.
-Detecting WebShells with LLMs presents unique challenges that differ from other code analysis tasks.
-WebShells often employ obfuscation and encryption techniques and are embedded in large codebases dominated by benign content (Liu & He, 2023).
-For example, the largest WebShell in our dataset spans 1,386,438 tokens, far exceeding the context window of most LLMs, which risks truncating critical malicious segments when processing entire source files (Wang et al., 2025; Ceka et al., 2024).
-In addition, in-context learning (ICL) struggles in this domain: the variability and obfuscation of WebShells complicate the selection of effective demonstrations, and these examples further occupy significant context space, reducing capacity for the target code (Yuan et al., 2024).
-While recent research has focused on increasing the context length of LLMs (Chen et al., 2023), studies suggest that the performance of an LLM tends to degrade with longer inputs, and the low processing speed may also become unacceptable for practical use (Ma et al., 2024; Fang et al., 2024).
-In this paper, we present two key contributions to advance the application of LLMs in WebShell detection.
-First, we systematically evaluate LLMs in the context of WebShell detection, comparing their performance with traditional state-of-the-art (SOTA) machine learning methods.
-Specifically, we evaluate seven closed-source and open-source LLMs of different sizes–including GPT-4 (Achiam et al., 2023), LLaMA 3.1 70B (Grattafiori et al., 2024), Qwen 2.5 Coder (14B/3B) (Yang et al., 2024), and Qwen 2.5 (3B/1.5B/0.5B) (Yang et al., 2024)–on a dataset containing 26.59K PHP scripts (4.93K WebShells and 21.66K benign samples).
-Our analysis reveals several key findings:
-  • Larger LLMs, such as GPT-4 and Qwen 2.5 Coder 14B, achieve near-perfect precision but struggle with recall (e.g., GPT-4’s recall is 85.98%).
-Their deep contextual analysis, while powerful, can be deceived by well-disguised malicious code, leading to false negatives.
-  • Smaller LLMs, such as Qwen 2.5 Coder 3B and Qwen 2.5 0.5B, exhibit high recall but suffer from low precision (e.g., Qwen 2.5 Coder 3B’s precision is 38.93%).
-This stems from their reliance on surface-level pattern matching, which often flags any use of sensitive functions without sufficient contextual validation.
-  • Randomly selected ICL demonstrations degrade LLM detection performance; examples selected based on semantic similarity to the source code do not yield significant improvements.
+The rapid growth of web applications and cloud services has expanded the attack surface of modern systems, making WebShells a persistent threat.
+A WebShell is a malicious script implanted on a web server that enables remote command execution, data exfiltration, and broader system compromise (Starov et al., 2016; Cui et al., 2018; Han, 2026).
+A Cisco Talos report (Talos, 2024) highlights the prevalence of such activity: web shells were observed in 35% of incidents in Q4 2024, up from 10% in the previous quarter.
+WebShells are particularly difficult to defend against because they are continually adapted—often via obfuscation or encryption—to evade conventional detection pipelines (Hannousse and Yahiouche, 2021).
+To counter this threat, prior work spans signature/heuristic defenses and learned detectors.
+Rule-based approaches are increasingly brittle given the diversity and rapid evolution of WebShell variants (Le et al., 2021; Jinping et al., 2020).
+Learning-based methods, including deep models (Pu et al., 2022), can capture richer patterns but typically require substantial labeled data and costly retraining—resources that are often scarce or sensitive in practice (Shang et al., 2024).
+Moreover, these models can suffer from catastrophic forgetting under continual updates and may generalize poorly to novel obfuscation or encryption strategies (Jinping et al., 2020; Zhang et al., 2025).
+Large language models (LLMs) have recently demonstrated strong performance on code-centric tasks, from generation to program understanding (Ma et al., 2024), and have been explored for security applications such as vulnerability detection (Liu and He, 2023; Wang et al., 2025; Han et al., 2026b).
+With appropriate prompting, LLMs can be adapted to new tasks without additional training (Nong et al., 2024; Trad and Chehab, 2025; Han et al., 2026a), and they can provide natural-language rationales that support analyst workflows (Ma et al., 2024).
+However, their effectiveness for WebShell detection remains insufficiently characterized.
+Applying LLMs to WebShell detection introduces challenges that are less pronounced in standard code analysis.
+WebShell payloads are often heavily obfuscated or encrypted and may be buried within predominantly benign code (Liu and He, 2023).
+Naively feeding an entire file to an LLM can therefore be ineffective: in our dataset, the longest WebShell contains 1,386,438 tokens, far beyond typical context windows, so truncation can easily remove the malicious core (Wang et al., 2025; Ceka et al., 2024).
+In addition, in-context learning (ICL) is fragile in this domain: the diversity of obfuscation patterns complicates demonstration selection, and demonstrations themselves consume context budget that would otherwise be allocated to the target code (Yuan et al., 2024).
+While longer-context LLMs are actively studied (Chen et al., 2023), prior work suggests that performance can degrade on longer inputs and latency may become prohibitive in practice (Ma et al., 2024; Fang et al., 2024).
+In this paper, we take a step toward making LLM-based WebShell detection both measurable and reliable.
+First, we conduct a systematic evaluation of LLMs for WebShell detection and compare them with representative learned detectors.
+Concretely, we evaluate seven closed- and open-source LLMs spanning a wide range of scales—including GPT-4 (Achiam et al., 2023), LLaMA 3.1 70B (Grattafiori et al., 2024), Qwen 2.5 Coder (14B/3B) (Yang et al., 2024), and Qwen 2.5 (3B/1.5B/0.5B) (Yang et al., 2024)—on a dataset of 26.59K PHP scripts (4.93K WebShells and 21.66K benign samples).
 For this comparison, we benchmark against several traditional methods, including GloVe+SVM (Petridis, 2024; Rigutini et al., 2024), CodeBERT+Random Forest (Alghamdi et al., 2022), and graph-based approaches such as GAT (Kang et al., 2023).
-The best-performing LLM, Qwen 2.5 Coder 14B, achieves an F1 score of 96.39%, although it still lags behind GAT-based methods, which achieve an F1 score of 98.87%.
-[^1] While traditional models such as GAT have demonstrated superior performance, they require extensive training and significant data collection efforts.
-In contrast, LLMs enable direct detection through prompts, leveraging their pre-trained knowledge with minimal task-specific resources to deliver competitive results.
-Second, we present the Behavioral Function-Aware Detection (BFAD) framework to address the identified challenges and improve the performance of LLMs to meet the requirements of downstream applications by achieving a more balanced trade-off between precision and recall.
-Our framework combines risk-filtering techniques with an enhanced ICL strategy that uses weighted demonstration selection to prioritize examples most closely related to key malicious behaviors.
-Experimental results show that our approach improves the average F1 score across all LLMs by 13.82%, with GPT-4 and Qwen 2.5 0.5B improving by 6.89% and 51.23%, respectively.
-In several cases, such as GPT-4, LLaMA 3.1 70B, Qwen 2.5 Coder 14B, and Qwen 2.5 Coder 3B, our approach enables performance that is competitive with or even superior to traditional methods.
-To the best of our knowledge, this is the first work to systematically analyze the feasibility and limitations of applying LLMs to WebShell detection.
+Our experiments yield three main observations:
+  • Scale shifts the error mode: larger LLMs (e.g., GPT-4, Qwen 2.5 Coder 14B) achieve very high precision but lower recall (e.g., GPT-4 recall 85.98%), whereas smaller models often trade precision for recall (e.g., Qwen 2.5 Coder 3B precision 38.93%).
+  • Naive ICL is unreliable: randomly selected demonstrations can degrade performance, and selecting demonstrations by semantic similarity alone yields limited improvements.
+  • LLMs do not close the gap by default: without task-specific adaptation, off-the-shelf LLM prompting remains behind strong learned detectors in overall F1.
+For example, the best-performing LLM baseline, Qwen 2.5 Coder 14B, achieves an F1 score of 96.39%, but still trails the GAT-based detector (98.87% F1).
+[^1] While traditional models such as GAT can achieve higher accuracy, they typically require task-specific training and sustained data collection.
+In contrast, LLMs can be deployed via prompting, leveraging pretraining with minimal task-specific supervision.
+Second, we propose Behavioral Function-Aware Detection (BFAD), a behavior-centric framework designed to make LLM inference effective under these constraints and to better balance precision and recall.
+BFAD couples risk-aware filtering with an enhanced ICL strategy that weights demonstrations by their alignment with discriminative malicious behaviors.
+Across all evaluated LLMs, BFAD improves F1 by 13.82% on average; GPT-4 and Qwen 2.5 0.5B improve by 6.89% and 51.23%, respectively.
+For several models (including GPT-4, LLaMA 3.1 70B, Qwen 2.5 Coder 14B, and Qwen 2.5 Coder 3B), BFAD yields performance competitive with—and in some cases surpassing—strong traditional baselines.
+Finally, we outline several forward-looking research directions suggested by our analysis, including (i) constructing privacy-preserving, large-scale synthetic benchmarks for stress-testing generalization, (ii) integrating graph-based behavioral representations with LLMs via multimodal alignment (e.g., graph encoders plus lightweight adapters), and (iii) designing agentic “fast–slow” detection pipelines with autonomous update loops to sustain performance under distribution shift.
+To the best of our knowledge, this is the first work to systematically characterize when and why LLMs succeed or fail on WebShell detection, and to provide a practical framework that closes much of the gap to learned state-of-the-art detectors.
+2 Related Work
+WebShell Detection Techniques.
+Early WebShell detection systems were predominantly rule-based, relying on signatures and handcrafted heuristics to flag malicious scripts (Le et al., 2021; Jinping et al., 2020).
+While effective against known families, such approaches are brittle under obfuscation and rapid variant evolution because they depend on predefined patterns (Hannousse and Yahiouche, 2021).
+Subsequent machine-learning (ML) methods introduced feature-based classification using textual or behavioral signals.
+For instance, Jinping et al.
+(2020) combined Random Forest and CNN models with N-gram and TF-IDF features, reporting strong performance on PHP WebShells but noting sensitivity to dataset balance and encrypted samples.
+Deep learning has further expanded the design space by leveraging semantic and structural representations, including transformer-based embeddings (e.g., CodeBERT) (Pu et al., 2022) and graph neural networks for code structure modeling.
+Zhang et al.
+(2025) proposed MMFDetect, which fuses CodeBERT-CL semantics with CNN-extracted visual features from RGB-mapped PHP code.
+Despite steady progress, supervised ML/DL detectors typically require substantial labeled data (often scarce in security settings) and can generalize poorly to heavily obfuscated or previously unseen WebShell variants, while incurring non-trivial training and maintenance costs (Shang et al., 2024; Jinping et al., 2020).
+LLMs for Code and Security Analysis.
+Large language models (LLMs) trained on massive code corpora have become a powerful interface for code understanding and generation, enabling applications ranging from synthesis (Ma et al., 2024) to vulnerability detection (Sun et al., 2024) and reliability assessment (Liu et al., 2024).
+In security contexts, LLMs are attractive because they support zero-shot and few-shot generalization via prompting and can provide natural-language rationales that improve analyst interpretability (Nong et al., 2024; Han et al., 2026c).
+For example, Ma et al.
+(2024) showed that LLMs can generate evasive WebShells with carefully designed prompts, and Sun et al.
+(2024) proposed retrieval-enhanced prompting for vulnerability reasoning.
+However, compared with these broader code-security applications, LLM-based WebShell detection remains relatively underexplored, especially for long, obfuscated scripts where malicious behavior may be sparsely embedded.
+Challenges in Applying LLMs to WebShell Detection.
+Directly applying LLMs to WebShell detection faces two practical bottlenecks.
+First, long WebShell files can exceed fixed context windows, causing truncation that may omit the truly malicious region (Ceka et al., 2024; Wang et al., 2025).
+Moreover, performance can degrade as input length grows; Fang et al.
+(2024) reported substantial accuracy drops for obfuscated long-context code.
+Techniques such as chunking and long-context model architectures (e.g., sparse attention) can partially alleviate this issue (Guo et al., 2023; Wang et al., 2024b), but may sacrifice global context and introduce fragmentation (Wang et al., 2025).
+Second, in-context learning (ICL) is itself context-expensive: demonstrations consume a significant fraction of the prompt budget (Min et al., 2022; Wang et al., 2025), and naive retrieval (random or whole-file semantic similarity) often fails to capture the behavioral structure of WebShells.
+While information-theoretic criteria have been explored for ICL selection (Liu and Wang, 2023), existing formulations primarily target natural language classification and do not directly address behavior-centric code security tasks.
+Taken together, the limitations of (i) rule-based rigidity, (ii) supervised ML/DL dependence on labeled data and weak OOD robustness, and (iii) LLM context and ICL constraints motivate BFAD.
+BFAD addresses long-context inputs via hybrid, behavior-anchored extraction and improves ICL reliability with WBFP, which retrieves demonstrations based on behavior-weighted, function-level profiles rather than generic similarity.
+3 Behavioral Function-Aware Detection Framework
+We present Behavioral Function-Aware Detection (BFAD), a framework that improves WebShell detection by (i) isolating behavior-relevant code snippets for LLM analysis and (ii) selecting in-context learning (ICL) demonstrations that better match the target file’s malicious behaviors.
+As illustrated in Figure 1, BFAD comprises three components: (a) a Critical Function Filter that maps PHP function calls to a behavior taxonomy and extracts behavior-relevant functions; (b) Context-Aware Code Extraction that constructs a compact yet informative code view under LLM context limits; and (c) Weighted Behavioral Function Profiling that ranks candidate demonstrations using a behavior-weighted similarity score computed from function-centric representations.
 Figure 1: Overview of the Behavioral Function-Aware Detection framework for WebShell detection.
 It consists of three components: (a) Critical Function Filter, which identifies PHP functions associated with malicious behavior; (b) Context-Aware Code Extraction, which isolates critical code regions to overcome LLM context limitations; and (c) Weighted Behavioral Function Profiling, which selects ICL demonstrations using a behavior-weighted similarity score.
-2 Behavioral Function-Aware Detection Framework
-We present the Behavioral Function-Aware Detection (BFAD) framework, a comprehensive solution designed to improve WebShell detection by identifying critical code segments and improving the quality of ICL demonstration selection.
-The architecture of BFAD, as shown in Figure 1, consists of three primary components: (a) Critical Function Filter, which is based on malicious behavior patterns and filters critical PHP function calls to identify key functions; (b) Context-Aware Code Extraction, which addresses the limitations of LLMs in handling long input sequences by selectively extracting critical code regions; and (c) Weighted Behavioral Function Profiling, which improves demonstration selection by calculating a weighted similarity score based on function-level profiling.
 3.1 Critical Function Filter
-WebShells typically rely on specific PHP functions that facilitate malicious actions such as code execution, data exfiltration, or obfuscation.
-However, these functions are often embedded in complex, obfuscated code, making detection difficult.
-To address this, we are developing a Critical Function Filter that classifies PHP functions into six different behavioral categories: Program Execution, Code Execution, Callback Functions, Network Communication, Information Gathering, and Obfuscation and Encryption.
-This taxonomy reflects the different roles these functions play in WebShells.
-Specifically, the Program Execution category includes functions such as exec and system, which execute system-level commands and are often exploited in WebShells to execute arbitrary payloads.
-Similarly, Code Execution functions such as eval and preg_replace interpret input as executable code, allowing attackers to inject arbitrary scripts.
-The Callback Functions category includes functions such as array_map and register_shutdown_function, which allow dynamic invocation of functions often used to obfuscate malicious code.
-In addition, Network Communication functions such as fsockopen and curl_init allow remote communication for data exfiltration and command-and-control operations.
-Information Gathering functions, such as phpinfo and getenv, are used by attackers to gather system details.
-Finally, Obfuscation and Encryption functions, such as base64_encode and openssl_encrypt, help disguise or encrypt malicious payloads to avoid detection.
-Our statistical analysis (detailed in Appendix C) shows that WebShell files use critical functions far more often than benign files.
-On average, WebShells contain 22.76 calls to critical functions, compared to only 0.74 in benign files, underscoring their behavioral complexity and distinctiveness.
-[^2] While this disparity in critical function usage is significant, it alone does not reliably distinguish WebShells from benign files, since legitimate scripts may also call these functions.
-Therefore, we use LLMs for deeper contextual analysis to improve detection accuracy.
+WebShells often implement malicious behaviors—e.g., command execution, data exfiltration, and payload obfuscation—through a small set of high-risk PHP functions.
+In practice, these calls are frequently buried in long, noisy, and obfuscated code, which dilutes LLM attention and complicates robust detection.
+BFAD therefore begins with a Critical Function Filter that organizes PHP functions into six behavior categories: Program Execution, Code Execution, Callback Functions, Network Communication, Information Gathering, and Obfuscation and Encryption.
+This taxonomy captures the distinct roles that functions play in typical WebShell workflows.
+Concretely, Program Execution includes functions such as exec and system that invoke system commands; Code Execution includes eval and preg_replace (with the /e modifier in legacy PHP) that can interpret strings as code; and Callback Functions includes array_map and register_shutdown_function that enable indirect invocation and are frequently used to hide control flow.
+Network Communication includes fsockopen and curl_init for remote communication (e.g., command-and-control or exfiltration); Information Gathering includes phpinfo and getenv for environment reconnaissance; and Obfuscation and Encryption includes base64_encode and openssl_encrypt for disguising or encrypting payloads.
+Our statistical analysis (Appendix B) indicates that WebShell files invoke these critical functions far more frequently than benign files: on average, WebShells contain 22.76 critical-function calls, compared to 0.74 in benign files.
+This gap suggests strong behavioral signals, but it is insufficient by itself because benign programs may legitimately use the same APIs.
+We therefore treat critical functions as anchors to locate behavior-relevant code and rely on LLMs to disambiguate benign versus malicious intent from context.
 3.2 Context-Aware Code Extraction
-Building on the Critical Function Filter, we introduce a Context-Aware Code Extraction strategy that identifies and extracts the critical code regions that indicate malicious behavior.
-These regions focus on the identified critical functions and their surrounding contexts, ensuring that the LLM focuses on the most relevant parts of the code.
+Building on the Critical Function Filter, BFAD performs Context-Aware Code Extraction to construct an LLM-friendly representation of each file.
+The key idea is to center extraction around critical function occurrences and include their surrounding context, so that the LLM receives the minimal code necessary to infer the intent of each call.
 The complete extraction procedure is formalized in Algorithm 1, which takes as input the source code $\mathcal{C}$, the list of critical functions $\mathcal{F}$, and the context window size $\tau$, and produces a set of extracted critical code regions $\mathcal{R}$.
-We reduce the input size by selectively extracting critical regions of code and merging overlapping segments while preserving behavioral specificity.
-However, this approach may inadvertently exclude certain global contextual information and increase the emphasis on critical functions, potentially leading to false positives when analyzing benign files that legitimately use such functions.
-To mitigate this, we append truncated, non-overlapping code segments when context length allows, ensuring that the model receives a balanced representation of local and global code context.
+We reduce input length by extracting windows around critical calls and merging overlapping windows, which preserves local behavioral evidence while avoiding redundant tokens.
+This operation can remove global context and over-emphasize critical functions, potentially increasing false positives for benign scripts that legitimately call high-risk APIs.
+To mitigate this, when the context budget allows we append additional truncated, non-overlapping segments from the remaining code, providing complementary global signals without exceeding the model’s context limit.
 Algorithm 1 Context-Aware Code Extraction
@@
 3.3 Weighted Behavioral Function Profiling
-Based on the extracted critical code regions, we propose Weighted Behavioral Function Profiling (WBFP), a method that computes a weighted similarity score to identify behaviorally similar examples for ICL effectively.
-WBFP assigns weights to each function type based on its prevalence and usage in WebShell versus benign files, quantified by three metrics: coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$).
-The coverage difference measures the proportion of files containing a specific function across the two datasets.
-The frequency ratio is the ratio of the average number of occurrences of the function per file in WebShell files to the average number per file in benign files.
-The usage ratio reflects the total number of function occurrences in WebShell files compared to those in benign files.
-These metrics are combined to calculate a discrimination score for each function type
+Given extracted regions, BFAD selects ICL demonstrations via Weighted Behavioral Function Profiling (WBFP).
+WBFP ranks candidate examples by a behavior-weighted similarity that emphasizes function types that best discriminate WebShells from benign files.
+Specifically, for each critical function type $f$, we compute three statistics from the WebShell and benign corpora: coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$).
+Coverage difference measures how often $f$ appears at least once in a file across the two corpora; frequency ratio measures the ratio of the average per-file call count; and usage ratio measures the ratio of the total call counts.
+We combine them into a discrimination score:
 $\text{Score}_{f}=(r_{c}\cdot\alpha)+(r_{f}\cdot\beta)+(r_{u}\cdot\gamma),$
-where $\alpha$, $\beta$, and $\gamma$ are empirically determined weights, set to 1 for balanced contribution in our experiments (see Section 3.2). We normalize the discrimination scores to weights:
+where $\alpha$, $\beta$, and $\gamma$ are scalar hyperparameters (set to 1 in our experiments for equal contribution; see Section 4.2). We normalize scores into weights:
 $w_{f}=\frac{\text{Score}_{f}}{\sum_{f^{\prime}\in\mathcal{F}}\text{Score}_{f^{\prime}}}.$
-WBFP then uses the embeddings $E(\cdot)$ generated by st-codesearch-distilroberta-base (Abi Akl, 2023; Al-Kaswan et al., 2023) to compute the similarity between the files $x$ and $y$.
-Let $\mathcal{F}$ denote all critical function types.
-For each function type $f\in\mathcal{F}$, we concatenate critical regions $R_{f}(x)$ and compute their embeddings:
-$\mathbf{e}_{f}(x)=E(\text{concat}_{f}(x)),$
-The semantic similarity between the files $x$ and $y$ for the function type $f$ is given by
+WBFP then computes similarity using embeddings $E(\cdot)$ from st-codesearch-distilroberta-base (Abi Akl, 2023; Al-Kaswan et al., 2023).
+Let $\mathcal{F}$ denote the set of critical function types.
+For each $f\in\mathcal{F}$, we concatenate the extracted regions associated with $f$ (denoted $R_{f}(x)$) and embed them as
+$\mathbf{e}_{f}(x)=E(\text{concat}_{f}(x)).$
+The function-type similarity between files $x$ and $y$ is computed by cosine similarity:
 $s_{f}(x,y)=\frac{\mathbf{e}_{f}(x)\cdot\mathbf{e}_{f}(y)}{\|\mathbf{e}_{f}(x)\|\|\mathbf{e}_{f}(y)\|}.$
@@
 $\text{Sim}(x,y)=\sum_{f\in\mathcal{F}}w_{f}\cdot s_{f}(x,y).$
-This weighted similarity prioritizes function types critical to WebShell behavior, reducing the impact of irrelevant semantic features.
-Using this score, WBFP ensures that ICL demonstrations accurately capture malicious patterns, improving detection performance.
+This formulation prioritizes function types that are more indicative of WebShell behavior and down-weights less informative semantics, yielding demonstrations that better match the target file’s behavioral profile.
 3.4 LLM-Based Detection Framework
-We integrate the BFAD framework into the LLM-based detection system, which combines the context-aware code extraction strategy and WBFP to optimize the use of LLMs for WebShell detection.
-By incorporating both critical code regions and global context, along with behaviorally relevant demonstrations, our framework enhances the ability of the LLM to accurately identify malicious code patterns.
-The input to the LLM consists of two main components: (a) a system directive that defines the model’s role as a cybersecurity expert, and (b) a user query that contains the extracted critical code segments and a selected ICL demonstration.
-To balance efficiency and performance, we limit the user query to one ICL demonstration, which reduces computational overhead while preserving sufficient context for reliable detection.
-Our prompt is described in detail in Appendix B.
+We integrate BFAD into an LLM-based detection pipeline that combines context-aware extraction with behavior-weighted demonstration selection.
+By providing behavior-relevant snippets together with complementary global context and a closely matched demonstration, the LLM is better positioned to distinguish malicious intent from benign uses of similar APIs.
+The LLM input consists of (a) a system directive specifying the cybersecurity-analysis role and (b) a user query containing the extracted code view and the selected ICL demonstration.
+To balance effectiveness and cost, we include a single demonstration per query.
+The full prompt template is provided in Appendix A.
 4 Experiment
 4.1 Dataset Overview
-We constructed a comprehensive dataset consisting of 21,665 benign PHP programs and 4,929 WebShells.
-The benign programs were obtained from established open-source PHP projects to ensure applicability to real-world scenarios.
-The WebShells were collected from public security repositories and augmented with synthetic obfuscation techniques to increase diversity.
-[^3] We acknowledge the potential for data leakage due to the pre-training of LLMs.
-However, the autoregressive nature of this process does not explicitly capture WebShell classification, although it may affect other code generation tasks.
-To further minimize any risk of leakage, we restricted our selection of PHP programs to projects updated between October 2024 and 2025, thus reducing overlap with LLM training corpora.
-Using GPT-4’s tokenizer, we analyzed the token lengths of both sample types.
-The WebShell samples had a maximum token length of 1,386,438 and an average of 30,856.60 tokens, compared to a maximum of 305,670 tokens and an average of 2,242.89 tokens for benign programs.
-These results indicate that WebShells are typically significantly longer than benign samples.
-A detailed summary of the dataset composition can be found in the Table 3 in the Appendix D.
+We construct a dataset of 26,594 PHP scripts, including 21,665 benign programs and 4,929 WebShells.
+Benign programs are sourced from established open-source PHP projects to reflect realistic development practices.
+WebShells are collected from public security repositories and further augmented with synthetic obfuscation to increase behavioral and lexical diversity.
+[^2] We acknowledge the risk of data leakage due to LLM pre-training.
+While pre-training does not explicitly optimize for WebShell classification, memorization of code snippets could still affect evaluation.
+To reduce overlap with potential training corpora, we restrict benign projects to those updated between October 2024 and 2025.
+To characterize the distributional shift and the challenge posed by long contexts, we measure sequence lengths using the GPT-4 tokenizer.
+WebShells are substantially longer than benign files, with a maximum of 1,386,438 tokens and an average of 30,856.60 tokens, compared to a maximum of 305,670 and an average of 2,242.89 tokens for benign programs.
+A detailed breakdown of the dataset is provided in Table 3 (Appendix C).
 4.2 Experiment Setup
 ICL Settings.
-We randomly selected 60% of the dataset to create a fixed, known demonstration library for ICL.
-Using this subset, we computed normalized scores for different function categories based on the WBFP method, giving equal weight to coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$) to profile functions according to their behavioral importance in distinguishing WebShells from benign programs.
-These scores, detailed in the Table 4 in the Appendix D, guided the selection of ICL demonstrations from the known sample library.
+We randomly sample 60% of the dataset to form a fixed demonstration library for ICL.
+Using this subset, we compute WBFP-based weights for each critical function type by combining coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$).
+The resulting normalized weights quantify each function type’s discriminative importance for separating WebShells from benign programs and are used to rank candidate demonstrations.
+The full weights are reported in Table 4 (Appendix C).
 Baseline Models, Hyperparameter Settings, and Evaluation Metrics.
-We compared our approach to several baselines: GloVe + SVM, CodeBERT + Random Forest, GCN, and GAT.
-For GloVe + SVM, we used pre-trained GloVe embeddings with a dimensionality of 300 and an SVM classifier with default parameters (Qi et al., 2018; ZENG et al., 2025).
-For CodeBERT + Random Forest, we used CodeBERT embeddings with a hidden dimension of 768 and a Random Forest classifier with default settings (Wang et al., 2024a).
-The GCN was trained with a learning rate of 0.001 over 120 epochs, with 3 hidden layers and a hidden dimension of 32 (Feng et al., 2024).
-The GAT was trained with a learning rate of 0.001 over 120 epochs, with 3 hidden layers, a hidden dimension of 8, and 8 attention heads (Feng et al., 2024).
-We evaluated the models using standard classification metrics: accuracy, precision, recall, and F1 score.
+We compare BFAD-enhanced LLMs against representative classical and graph-based detectors: GloVe+SVM, CodeBERT+Random Forest, GCN, and GAT.
+For GloVe+SVM, we use 300-dimensional pre-trained GloVe embeddings and an SVM classifier with default hyperparameters (Qi et al., 2018; ZENG et al., 2025).
+For CodeBERT+Random Forest, we use CodeBERT embeddings (768 dimensions) with a Random Forest classifier using default settings (Wang et al., 2024a).
+For the graph models, we follow prior work (Feng et al., 2024): both GCN and GAT are trained for 120 epochs with learning rate 0.001; GCN uses 3 hidden layers with hidden size 32, and GAT uses 3 hidden layers with hidden size 8 and 8 attention heads.
+We report standard classification metrics: accuracy, precision, recall, and F1.
 5 Results and Analysis
-In this section, we systematically evaluate the performance of LLMs of different scales for WebShell detection and assess the improvements provided by our proposed BFAD framework.
-Our analysis addresses three research questions (RQs) to explore both the baseline LLM capabilities and the effectiveness of the BFAD components:
+This section evaluates LLMs of different scales for WebShell detection and quantifies the gains delivered by our proposed BFAD framework.
+We structure the analysis around three research questions (RQs), covering both out-of-the-box LLM behavior and the contribution of individual BFAD components:
   • RQ1: How do large and small-scale LLMs perform in WebShell detection compared to traditional ML and DL methods, and how does BFAD improve their effectiveness?
@@
 5.1 Performance Evaluation of LLMs and BFAD Enhancements (RQ1)
-To answer our first research question, we evaluated seven LLMs against traditional ML and DL baselines.
-The LLMs included large-scale models (GPT-4, LLaMA-3.1-70B, Qwen-2.5-Coder-14B) and small-scale models (Qwen-2.5-Coder-3B, Qwen-2.5-3B, Qwen-2.5-1.5B, Qwen-2.5-0.5B).
-Baselines consisted of sequence-based methods (Glove+SVM, CodeBERT+RF) and graph-based methods (GCN, GAT).
-The detailed performance comparison is presented in Table 1.
+To address RQ1, we compare seven LLMs with representative classical detectors.
+The LLMs include large-scale models (GPT-4, LLaMA-3.1-70B, Qwen-2.5-Coder-14B) and smaller models (Qwen-2.5-Coder-3B, Qwen-2.5-3B, Qwen-2.5-1.5B, Qwen-2.5-0.5B).
+Baselines include sequence-based methods (GloVe+SVM, CodeBERT+RF) and graph-based methods (GCN, GAT).
+Table 1 reports all metrics.
 Table 1: Performance Comparison of BFAD-Enhanced Models Against Baselines.
@@
 Performance Summary
-Our experimental results yield two core findings.
-First, vanilla LLMs exhibit a stark, size-dependent trade-off between precision and recall, with all models lagging behind the state-of-the-art GAT baseline.
-For instance, GPT-4 achieves 100% precision but only 85.98% recall, while the small Qwen-2.5-0.5B model reaches 100% recall but a mere 18.65% precision.
-This indicates that out-of-the-box LLMs are not well-calibrated for the nuanced task of WebShell detection, making targeted intervention necessary for effective deployment.
-Second, the BFAD framework dramatically improves performance across all models, elevating large LLMs to surpass the GAT baseline and making small models viable competitors.
-With BFAD, GPT-4’s F1 score increases by 6.89 points to 99.35%, exceeding the GAT baseline.
-The most significant gain is seen in Qwen-2.5-0.5B, whose F1 score improves by a remarkable 51.23 points.
-This demonstrates that by strategically guiding the model’s focus on behaviorally relevant code, BFAD effectively bridges the performance gap, unlocking the potential of LLMs for this task without resource-intensive retraining.
+Our results highlight two key observations.
+First, vanilla LLM prompting exhibits a pronounced, scale-dependent precision–recall trade-off and remains below the strongest graph baseline (GAT).
+For example, GPT-4 attains 100% precision but only 85.98% recall, whereas Qwen-2.5-0.5B achieves 100% recall with only 18.65% precision.
+This asymmetry suggests that off-the-shelf LLMs are poorly calibrated for WebShell detection, motivating behavior-aware intervention.
+Second, BFAD improves every evaluated LLM, elevating large models beyond the GAT baseline and making small models substantially more competitive.
+With BFAD, GPT-4 improves by +6.89 F1 to 99.35%, surpassing GAT.
+The largest relative gain occurs for Qwen-2.5-0.5B (+51.23 F1), indicating that explicitly steering attention toward behavior-relevant code and demonstrations can unlock strong performance without task-specific fine-tuning.
 Analysis of Model Behavior: Misclassification Patterns
-The performance disparities observed in our results appear to stem from fundamental differences in the models’ analytical reasoning.
-The precision-recall imbalance in baseline LLMs is a direct reflection of their reasoning strategy.
-Small models like Qwen-2.5-0.5B tend to rely on surface-level heuristics; they identify keywords like base64_decode and flag them as malicious without assessing the surrounding code’s intent.
-This leads to high recall but generates numerous false positives when such functions are used legitimately (e.g., in a cryptography library).
-In contrast, large models like LLaMA-3.1-70B employ deep contextual reasoning, correctly identifying legitimate use cases and thus achieving high precision.
-However, this sophistication can be a double-edged sword, as these models can be misled by well-disguised WebShells, rationalizing away malicious indicators and causing false negatives that reduce recall.
-Our proposed BFAD mitigates these issues by guiding smaller models to better understand function-level behaviors and helping larger models focus on critical functions.
-As a result, it significantly improves detection performance across scales.
+The above gap is consistent with qualitative misclassification patterns.
+Smaller models (e.g., Qwen-2.5-0.5B) often rely on surface heuristics: they over-trigger on high-risk API names such as base64_decode and ignore surrounding intent, yielding high recall but many false positives when such APIs appear in benign utilities (e.g., cryptography libraries).
+In contrast, larger models (e.g., LLaMA-3.1-70B) more frequently leverage broader context and therefore maintain high precision, but can be distracted by plausible-looking surrounding logic and miss carefully disguised WebShells, producing false negatives.
+BFAD mitigates both failure modes by anchoring analysis on behavior-critical calls while still preserving enough local (and partial global) context for intent disambiguation.
 Analysis of Model Behavior: Counterintuitive Scaling
-We observed a counterintuitive scaling effect where the Qwen-2.5-1.5B model, despite outperforming the 0.5B variant in baseline tests, performed worse with BFAD.
-We attribute this to differing sensitivities to ICL across scales.
-The smallest models, like the 0.5B variant, are highly susceptible to label bias and tend to mimic the ICL demonstration’s label without deep reasoning (Fei et al., 2023).
-Because BFAD provides accurately matched demonstrations via WBFP, the 0.5B model benefits disproportionately from this behavior.
-In contrast, the 1.5B model begins to exhibit more independent reasoning based on its pre-training knowledge.
-This emergent but still imperfect reasoning can conflict with the ICL example’s guidance, degrading performance.
-This aligns with prior findings that LLMs sometimes rely more on pre-training than on provided examples (Peng et al., 2025).
-Therefore, the 0.5B model’s superior performance with BFAD does not indicate greater intelligence but is an artifact of its strong alignment with the high-quality examples provided.
+We observe a counterintuitive scaling pattern: Qwen-2.5-1.5B, while stronger than the 0.5B variant under vanilla prompting, improves less under BFAD and remains worse than Qwen-2.5-0.5B with BFAD.
+A plausible explanation is scale-dependent sensitivity to ICL.
+Very small models are more prone to label bias and may over-imitate the demonstration label rather than perform independent reasoning (Fei et al., 2023); when WBFP retrieves well-matched demonstrations, this bias can translate into large gains.
+As model size increases, the model may rely more on pre-training priors and partially resist demonstration guidance, which can introduce conflicts and reduce the net benefit (Peng et al., 2025).
+Thus, the strong BFAD performance of the 0.5B model should be interpreted as a demonstration-following effect rather than superior underlying reasoning.
 5.2 The Effectiveness of Context-Aware Code Extraction (RQ2)
-We evaluated the effectiveness of Context-Aware Code Extraction using two models: GPT-4, a large-scale model, and Qwen-2.5-3B, a smaller-scale model.
-Three configurations were compared: (1) predictions based on the full source code, (2) predictions using only extracted critical regions, and (3) a hybrid approach combining critical regions with truncated source code.
-Results are reported in Tables 5 and 6 in Appendix E.
+To answer RQ2, we evaluate Context-Aware Code Extraction on a large model (GPT-4) and a smaller model (Qwen-2.5-3B).
+We compare three input configurations: (1) the full source code, (2) extracted critical regions only, and (3) a hybrid input that concatenates critical regions with additional truncated code segments.
+Results are reported in Tables 5 and 6 (Appendix D).
 Impact of Critical Regions
-For the smaller model, Qwen-2.5-3B, critical regions significantly improve performance over the full source baseline.
-At $\tau=100$, the F1 score increases from 84.37% to 90.91% (+6.54%), with precision increasing from 78.03% to 86.71% (+8.68%) and recall increasing from 91.84% to 95.54% (+3.70%).
-This improvement validates the focus on behaviorally relevant code snippets, which reduces irrelevant context and sharpens the focus of the model.
-However, as the context length ($\tau$) increases, performance decreases - F1 drops to 88.17% at $\tau=300$, likely due to the model’s limited ability to handle extended context, which introduces noise that degrades accuracy.
-For the larger model, GPT-4, critical regions increase recall but slightly decrease precision.
-At $\tau=300$, recall improves from 85.98% to 96.18% (+10.20%), while precision drops from 100.00% to 98.69% (-1.31%).
-The F1 score increases from 92.46% to 97.42% (+4.96%), indicating that GPT-4 effectively uses localized behavioral cues to improve recall.
-However, the reduced global context may introduce small biases, leading to a precision trade-off, although the overall performance remains strong due to the model’s greater capacity.
+For Qwen-2.5-3B, focusing on critical regions substantially improves performance over full-code prompting.
+At $\tau=100$, F1 increases from 84.37% to 90.91% (+6.54), with precision increasing from 78.03% to 86.71% (+8.68) and recall increasing from 91.84% to 95.54% (+3.70).
+This supports the hypothesis that removing irrelevant context sharpens behavioral evidence for smaller models.
+However, increasing $\tau$ can re-introduce noise: at $\tau=300$, F1 drops to 88.17%, consistent with limited robustness to longer, noisier contexts.
+For GPT-4, critical regions mainly improve recall with a minor precision trade-off.
+At $\tau=300$, recall rises from 85.98% to 96.18% (+10.20), while precision decreases from 100.00% to 98.69% (-1.31), resulting in an F1 increase from 92.46% to 97.42% (+4.96).
+This indicates that localized behavioral cues are particularly helpful for recovering missed attacks, while removing global context can slightly increase false positives.
 Balancing Precision and Recall with the Hybrid Strategy
-The hybrid strategy improves model performance by effectively balancing precision and recall.
-For Qwen-2.5-3B, it increases precision over using critical regions alone, from 86.71% to 89.02% (+2.31%) for $\tau=100$ and from 82.32% to 85.55% (+3.23%) for $\tau=300$.
-Although recall decreases slightly, the F1 score increases to 89.70% at $\tau=300$ (+1.53% from 88.17%), indicating that the hybrid approach reduces noise in longer contexts and supports smaller models by maintaining focus on critical regions while integrating valuable global context.
-For GPT-4, the strategy improves recall without compromising precision: at $\tau=300$, recall increases from 85.98% (using the source code) to 96.82% (+10.84%), while precision remains steady at 100.00%, yielding an F1 score of 98.38%–a 5.92% gain over the source code and a 0.96% improvement over critical regions alone.
+The hybrid strategy improves the precision–recall balance by re-introducing limited global context.
+For Qwen-2.5-3B, it increases precision relative to critical-regions-only (86.71% $\rightarrow$ 89.02% at $\tau=100$; 82.32% $\rightarrow$ 85.55% at $\tau=300$).
+Although recall decreases slightly, the hybrid input yields a higher F1 at $\tau=300$ (89.70%, +1.53 over 88.17), suggesting that carefully chosen global snippets reduce false positives when longer windows add noise.
+For GPT-4, the hybrid input improves recall without sacrificing precision: at $\tau=300$, recall increases from 85.98% (full code) to 96.82% (+10.84) while precision remains 100.00%, yielding an F1 of 98.38% (+5.92 over full code and +0.96 over critical-regions-only).
 5.3 The Effectiveness of WBFP for In-Context Learning (RQ3)
-We evaluated the effectiveness of WBFP for ICL demonstration selection using Qwen-2.5-3B and GPT-4.
-This evaluation builds on the optimal hybrid strategy from RQ2.
-Four demonstration selection strategies were compared: Random Selection (Random), Source Code Semantic Similarity (SC-Sim), WBFP with Equal Weights (WBFP-Eq), and WBFP with Function-Level Weights (WBFP-Wt).
-Results are detailed in Tables 7 and 8 in Appendix E.
+To answer RQ3, we evaluate WBFP for ICL demonstration selection on Qwen-2.5-3B and GPT-4, using the best-performing hybrid input from RQ2.
+We compare four retrieval strategies: Random selection, source-code semantic similarity (SC-Sim), WBFP with equal weights (WBFP-Eq), and WBFP with function-level weights (WBFP-Wt).
+Results are reported in Tables 7 and 8 (Appendix D).
 Limitations of Random and Semantic Similarity-Based Selection
-Random selection makes ICL performance much worse by adding irrelevant examples, which makes both models much less effective.
-For Qwen-2.5-3B, the F1 score drops to 60.83% under the hybrid strategy with a context length of $\tau=100$, which is a 30.14% reduction from the no-ICL baseline of 90.97%.
-This is mainly because there is a large drop in precision from 89.02% to 46.33%, although recall remains strong at 88.53%.
-Similarly, for GPT-4 with $\tau=300$, the F1 score drops to 76.22%, a 22.16% decrease from the No-ICL baseline of 98.38%, driven by a drop in precision from 100.00% to 65.58%.
-These results indicate that Random Selection does not provide contextually relevant demonstrations, making it ineffective for WebShell detection.
-The SC-Sim approach, which relies on semantic similarity calculated over entire source code samples, also doesn’t work well.
-For Qwen-2.5-3B, SC-Sim achieves an F1 score of 84.36%, a notable decrease from the No-ICL baseline of 90.97%, with a precision of 78.57% and a recall of 91.08%.
-For GPT-4, it achieves an F1 score of 96.32%, with perfect precision, but a reduced recall of 92.90% compared to the No-ICL baseline’s 96.82%.
-This limited performance is likely due to the dominance of behaviorally irrelevant code segments in the similarity calculation, which dilutes the focus on critical behavioral patterns essential for accurate WebShell identification.
+Random demonstrations substantially degrade ICL performance, indicating that irrelevant examples introduce strong prompt noise.
+For Qwen-2.5-3B ($\tau=100$), F1 drops to 60.83% (from 90.97% without ICL), driven primarily by a precision collapse (89.02% $\rightarrow$ 46.33%) while recall remains high (88.53%).
+For GPT-4 ($\tau=300$), F1 decreases to 76.22% (from 98.38%), largely due to reduced precision (100.00% $\rightarrow$ 65.58%).
+These results show that naive ICL can be harmful without behaviorally relevant retrieval.
+SC-Sim, which computes semantic similarity over entire source files, is also suboptimal.
+For Qwen-2.5-3B, SC-Sim yields 84.36% F1 (vs.
+90.97% without ICL), with precision 78.57% and recall 91.08%.
+For GPT-4, SC-Sim reaches 96.32% F1 with perfect precision but reduced recall (92.90% vs.
+96.82%).
+We hypothesize this occurs because whole-file similarity is dominated by behavior-irrelevant regions, which dilutes the behavioral cues needed for effective WebShell retrieval.
 Superiority of WBFP in Demonstration Selection
-For both models, WBFP-Wt consistently outperforms in accuracy, precision, recall, and F1 score, demonstrating its robustness and adaptability for improving ICL in WebShell detection tasks.
-Specifically, for Qwen-2.5-3B, WBFP-Wt achieves an F1 score of 93.69%, which is 9.33% higher than SC-Sim and 1.28% higher than WBFP-Eq.
-This is because WBFP-Wt achieves a precision of 88.64% (compared to 78.57% for SC-Sim and 86.39% for WBFP-Eq) and a near-perfect recall of 99.36%.
-By focusing on the important parts of WebShell behavior, WBFP-Wt makes up for the fact that the smaller model doesn’t understand as much.
-This results in demonstrations that closely match the desired behavioral profiles, improving both precision and recall.
-For GPT-4, WBFP-Wt achieves the highest F1 score of 99.35%, which is 3.03% higher than SC-Sim and 0.34% higher than WBFP-Eq .
-While precision remains at 100.00% across all WBFP variants, WBFP-Wt increases recall to 98.71%, compared to 92.90% for SC-Sim and 98.03% for WBFP-Eq.
-This improvement highlights WBFP-Wt’s ability to leverage GPT-4’s advanced understanding of context, matching selected demonstrations to the behavioral characteristics of the target sample to optimize recall without sacrificing precision.
-5 Related Work
-WebShell Detection Techniques Early efforts in WebShell detection relied on rule-based methods that used signature matching or heuristics to identify malicious scripts (Le et al., 2021; Jinping et al., 2020).
-For example, Le et al.
-(2021) proposed H-DLPMWD, a hybrid approach that combines pattern matching with a CNN to detect ASP.NET WebShells, achieving 98.49% accuracy by using Yara-based filtering and opcode indexing.
-However, such methods struggle against obfuscated or novel WebShell variants due to their reliance on predefined patterns (Hannousse & Yahiouche, 2021).
-Machine learning (ML) techniques have advanced this landscape by extracting features from code text or runtime behavior.
-Jinping et al.
-(2020) introduced a mixed-model approach using Random Forest and CNNs with N-gram and TF-IDF features that achieved 97% accuracy on PHP WebShells, but it requires balanced datasets and struggles with encrypted samples.
-Deep learning (DL) has further improved adaptability, with models such as CodeBERT for semantic analysis (Pu et al., 2022) and Graph Attention Networks (GAT) for structural insights (Zhang et al., 2025).
-(Zhang et al., 2025) proposed MMFDetect, which fuses CodeBERT-CL semantics with CNN-extracted visual features from RGB-mapped PHP code, achieving 99.47% accuracy on evasive WebShells.
-Despite these gains, ML and DL methods require extensive labeled data-a scarcity in cybersecurity-and exhibit limited generalization to obfuscated threats, along with high computational cost (Shang et al., 2024; Jinping et al., 2020).
-LLMs in Code Analysis LLMs have revolutionized code-related tasks by leveraging large pre-training corpora for applications such as code generation (Ma et al., 2024), vulnerability detection (Sun et al., 2024), and program reliability assessment (Liu et al., 2024).
-Ma et al.
-(2024) demonstrated the ability of LLMs to generate evasive WebShells using hybrid prompts, highlighting their code synthesis potential.
-Sun et al.
-(2024) introduced LLM4Vuln, which enhances vulnerability reasoning through knowledge retrieval and has achieved practical success in Solidity audits.
-These models excel at zero-shot and few-shot learning via prompt engineering (Nong et al., 2024), providing interpretability critical for security contexts (Ma et al., 2024).
-However, their application to WebShell detection remains underexplored, with previous studies focusing on generation or general vulnerabilities rather than detection of stealthy, context-heavy WebShells.
-Challenges of LLMs for WebShell Detection Using LLMs for WebShell detection reveals critical bottlenecks.
-The fixed context window truncates large WebShells, potentially missing malicious segments embedded in benign code (Ceka et al., 2024; Wang et al., 2025).
-Fang et al.
-(2024) found that LLM performance degrades with longer inputs, with GPT-4’s accuracy dropping to 87% on obfuscated JavaScript.
-Solutions such as chunking or sparse attention (e.g., LongCoder (Guo et al., 2023), SparseCoder (Wang et al., 2024b)) mitigate this, but often lose global context (Wang et al., 2025).
-In-context learning (ICL), a cornerstone of LLM adaptability, falters because demonstrations consume context space (Min et al., 2022; Wang et al., 2025), and random or semantic similarity-based selections fail to capture WebShell-specific behaviors.
-Liu & Wang (2023) proposed maximum information gain for ICL, but its focus on text classification limits its applicability to code security.
-These gaps between rule-based rigidity, ML and DL data dependency, and LLM context and ICL limitations motivate our BFAD framework.
-BFAD overcomes context constraints with a hybrid extraction strategy that preserves critical regions and global cues, outperforming naive LLM applications.
-In addition, our WBFP enhances ICL by prioritizing behaviorally relevant demonstrations, outperforming generic similarity-based selection methods.
+WBFP improves ICL reliably, and WBFP-Wt is consistently the strongest variant.
+For Qwen-2.5-3B, WBFP-Wt achieves 93.69% F1, improving over SC-Sim by +9.33 and over WBFP-Eq by +1.28.
+The gain comes from both higher precision (88.64% vs.
+78.57% for SC-Sim) and near-perfect recall (99.36%).
+These results suggest that behavior-weighted, function-centric retrieval is particularly beneficial for smaller models whose general code reasoning is limited.
+For GPT-4, WBFP-Wt achieves 99.35% F1 (+3.03 over SC-Sim and +0.34 over WBFP-Eq).
+Precision remains 100.00% across WBFP variants, while WBFP-Wt improves recall to 98.71% (vs.
+92.90% for SC-Sim and 98.03% for WBFP-Eq).
+This indicates that behavior-aware retrieval primarily helps large models recover hard positives without sacrificing precision.
+6 Limitations and Future Work
+In this section, we outline several concrete directions for future work—including synthetic benchmark construction, graph–LLM integration, and agentic detection pipelines—with the goal of making our ideas actionable and easy to build upon.
+6.1 Dataset Generalization
+A central challenge in WebShell research is the limited availability of diverse, real-world data.
+As in prior work, we primarily rely on public datasets, which may not fully reflect emerging, targeted, or highly obfuscated threats, and therefore limits our ability to rigorously assess generalization.
+An important direction is to use LLMs to construct large-scale synthetic benchmarks that are both privacy-preserving and diverse, enabling stress tests across a broader spectrum of attack vectors.
+6.2 Model and System Robustness
+Our framework is primarily static, and thus remains susceptible to advanced obfuscation and previously unseen (e.g., zero-day) behaviors.
+Future work should improve both robustness and adaptability.
+To mitigate evasive variants, one could incorporate dynamic analysis, e.g., by extracting runtime traces from sandboxed execution (Zhao et al., 2024; Han, 2026).
+To strengthen semantic understanding, domain-specific fine-tuning of code-oriented LLMs on curated WebShell corpora is another promising direction.
+Beyond standard fine-tuning, we propose a concrete next step that frames WebShell detection as a multimodal fusion problem: prior work suggests that graph representations can capture program behavior (e.g., function-call structure) more faithfully than pure sequence models (Han, 2026).
+Building on this observation, future work could (i) train a dedicated graph encoder to map code into behavior-centric embeddings (instead of relying solely on the LLM’s text embeddings), and then (ii) learn an adapter that aligns the graph embeddings with natural-language instructions and the LLM’s latent space (in the spirit of adapter-based alignment in VLMs and emerging graph-LLM systems).
+We believe this design can combine the strengths of both worlds: robust behavioral feature extraction from graphs, and the LLM’s pretrained discrimination and interpretability.
+For deployment, a practical direction is a multi-model agentic system that orchestrates specialized components.
+These components naturally fit a multi-stage “fast–slow” pipeline: a lightweight model performs high-recall triage, followed by a larger model that conducts deeper analysis on flagged samples.
+Finally, to sustain performance under distribution shift, the pipeline could be augmented with an autonomous update loop that discovers new patterns and refreshes detection models over time.
 7 Conclusion
-This paper systematically demonstrates that baseline LLMs are suboptimal for WebShell detection, exhibiting a precision-recall trade-off that underperforms state-of-the-art methods.
-We propose the Behavioral Function-Aware Detection (BFAD) framework, which overcomes these limitations by focusing analysis on critical code segments and optimizing in-context learning.
-BFAD enables our LLM-based approach to surpass traditional benchmarks.
-Appendix A Limitations and Future Work
-Our research also highlights several limitations that present important directions for future investigation in this emerging field.
-A.1 Dataset Generalization
-A fundamental challenge in WebShell research is the scarcity of diverse, real-world data.
-Our reliance on public datasets, like that of prior work, limits the evaluation of model generalization against novel, sophisticated threats.
-To address this, future work should focus on using LLMs to generate large-scale synthetic benchmarks.
-A large-scale, LLM-generated dataset would provide a privacy-preserving and diverse alternative for stress-testing model robustness against a wider range of attack vectors.
-A.2 Model and System Robustness
-Our framework’s static analysis approach remains vulnerable to advanced obfuscation and zero-day attacks.
-Future work should therefore focus on enhancing model robustness and system adaptability.
-To counter evasion, models could integrate dynamic analysis by examining runtime function call traces from sandboxed execution.
-To deepen semantic understanding, domain-specific fine-tuning of code-centric LLMs on curated WebShell datasets is a crucial next step.
-For practical deployment, these enhanced models could be integrated into a multi-stage ”fast and slow” system, using a smaller model for rapid filtering and a larger one for in-depth analysis of high-risk samples.
-To ensure long-term resilience, this pipeline could be governed by an autonomous agent that continuously identifies new attack patterns and automatically updates the detection models.
-Appendix B Prompt Details
+This paper provides a systematic study of LLMs for WebShell detection and shows that off-the-shelf prompting is not sufficient: across model scales, vanilla LLMs exhibit a pronounced precision–recall imbalance and fall short of strong supervised baselines.
+To bridge this gap, we introduce Behavioral Function-Aware Detection (BFAD), which adapts LLM inference to WebShell behaviors through (i) critical-function anchoring, (ii) context-aware extraction that fits long scripts into a limited prompt budget, and (iii) behavior-weighted demonstration retrieval via WBFP.
+Empirically, BFAD consistently improves all evaluated LLMs and enables several models to match or surpass state-of-the-art detectors.
+Overall, our findings clarify both the failure modes and the promise of LLM-based WebShell detection, and suggest that behavior-centric prompting and retrieval are key to making LLMs reliable in security-critical code analysis.
+Appendix A Prompt Details
 Prompt for WebShell Detection System Prompt: You are tasked with analyzing PHP scripts.
@@
 [Critical Code] [Source Code] [Examples] Output:
-Appendix C Critical Function Details
+Appendix B Critical Function Details
 Table 2: Statistics of Critical Functions in WebShell and Benign Programs.
@@
 Avg. Occurrences per File 22.76 0.74
-Appendix D Dataset Details
+Appendix C Dataset Details
 Table 3: Dataset Composition, Distribution, and Sources.
@@
 Obfuscation and Encryption 0.1702
-Appendix E Results
+Appendix D Results
 Table 5: Performance of Context-Aware Code Extraction with Qwen-2.5-3B with Different Context Lengths and Strategies.